### v3

# Recursive Neyman Algorithm for Optimum Sample Allocation under Box Constraints on Sample Sizes in Strata

Abstract
The optimum sample allocation in stratified sampling is one of the basic issues of survey methodology. It is a procedure of dividing the overall sample size into strata sample sizes in such a way that for given sampling designs in strata the variance of the stratified $\pi$ estimator of the population total (or mean) for a given study variable assumes its minimum. In this work, we consider the optimum allocation of a sample, under lower and upper bounds imposed jointly on sample sizes in strata. We are concerned with the variance function of some generic form that, in particular, covers the case of the simple random sampling without replacement in strata. The goal of this paper is twofold. First, we establish (using the Karush-Kuhn-Tucker conditions) a generic form of the optimal solution, the so-called optimality conditions. Second, based on the established optimality conditions, we derive an efficient recursive algorithm, named RNABOX, which solves the allocation problem under study. The RNABOX can be viewed as a generalization of the classical recursive Neyman allocation algorithm, a popular tool for optimum allocation when only upper bounds are imposed on sample strata-sizes. We implement RNABOX in R as a part of our package stratallo which is available from the Comprehensive R Archive Network (CRAN) repository.
Key Words: Stratified sampling; Optimum sample allocation; Optimum allocation; Optimum allocation under box constraints; Neyman allocation; Recursive Neyman algorithm.
1 Introduction
Let us consider a finite population $U$ of size $N$. Suppose the parameter of interest is the population total $t$ of a variable $y$ in $U$, i.e. $t=\sum_{k\in U}\,y_{k}$, where $y_{k}$ denotes the value of $y$ for population element $k\in U$. To estimate $t$, we consider the stratified sampling with the $\pi$ estimator. Under this well-known sampling technique, population $U$ is stratified, i.e. $U=\bigcup_{h\in\mathcal{H}}\,U_{h}$, where $U_{h},\,h\in\mathcal{H}$, called strata, are disjoint and non-empty, and $\mathcal{H}$ denotes a finite set of strata labels. The size of stratum $U_{h}$ is denoted $N_{h},\,h\in\mathcal{H}$ and clearly $\sum_{h\in\mathcal{H}}N_{h}=N$. Probability samples $s_{h}\subseteq U_{h}$ of size $n_{h}\leq N_{h},\,h\in\mathcal{H}$, are selected independently from each stratum according to chosen sampling designs which are often of the same type across strata. The resulting total sample is of size $n=\sum_{h\in\mathcal{H}}\,n_{h}\leq N$. It is well known that the stratified $\pi$ estimator $\hat{t}_{\pi}$ of $t$ and its variance are expressed in terms of the first and second order inclusion probabilities (see, e.g. Särndal et al., 1992, Result 3.7.1, p. 102). In particular, for several important sampling designs
$Var(\hat{t}_{\pi})=\sum_{h\in\mathcal{H}}\,\tfrac{A_{h}^{2}}{n_{h}}-B,$ (1.1)
where $A_{h}>0,\,B$ do not depend on $n_{h},\,h\in\mathcal{H}$. Among the most basic and common sampling designs that give rise to the variance of the form (1.1) is the simple random sampling without replacement in strata (abbreviated STSI). In this case, the stratified $\pi$ estimator of $t$ assumes the form
$\hat{t}_{\pi}=\sum_{h\in\mathcal{H}}\,\tfrac{N_{h}}{n_{h}}\sum_{k\in s_{h}}y_{k},$ (1.2)
which yields in (1.1): $A_{h}=N_{h}S_{h}$, where $S_{h}$ denotes stratum standard deviation of study variable $y$, $h\in\mathcal{H}$, and $B=\sum_{h\in\mathcal{H}}\,N_{h}S_{h}^{2}$ (see, e.g. Särndal et al., 1992, Result 3.7.2, p. 103).
The classical problem of optimum sample allocation is formulated as the determination of the allocation vector $\mathbf{n}=(n_{h},\,h\in\mathcal{H})$ that minimizes the variance (1.1), subject to $\sum_{h\in\mathcal{H}}\,n_{h}=n$, for a given $n\leq N$ (see, e.g. Särndal et al., 1992, Section 3.7.3, p. 104). In this paper, we are interested in the classical optimum sample allocation problem with additional two-sided constraints imposed on sample sizes in strata. We phrase this problem in the language of mathematical optimization as Problem 1.1.
Problem 1.1.
Given a finite set $\mathcal{H}\neq\emptyset$ and numbers $A_{h}>0,\,m_{h},\,M_{h},\,n$, such that $0<m_{h}<M_{h}\leq N_{h},\,h\in\mathcal{H}$ and $\sum_{h\in\mathcal{H}}\,m_{h}\leq n\leq\sum_{h\in\mathcal{H}}\,M_{h}$,
$\underset{\mathbf{x}\,=\,(x_{h},\,h\,\in\,\mathcal{H})\,\in\,{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}}{\mathrm{minimize\leavevmode\nobreak\ \,}}$ $\quad\sum_{h\in\mathcal{H}}\tfrac{A_{h}^{2}}{x_{h}}$ (1.3)
$\mathrm{subject\leavevmode\nobreak\ to\quad\,\,\,}$ $\quad\sum_{h\in\mathcal{H}}x_{h}=n$ (1.4)
$\quad m_{h}\leq x_{h}\leq M_{h},\mathrlap{\qquad h\in\mathcal{H}.}$ (1.5)
To emphasize the fact that the optimal solution to Problem 1.1 may not be an integer one, we denote the optimization variable by $\mathbf{x}$, not by $\mathbf{n}$. The assumptions about $n,\,m_{h},\,M_{h},\,h\in\mathcal{H}$, ensure that Problem 1.1 is feasible.
The upper bounds imposed on $x_{h},\,h\in\mathcal{H}$, are natural since for instance the solution with $x_{h}>N_{h}$ for some $h\in\mathcal{H}$ is impossible. The lower bounds are necessary e.g. for estimation of population strata variances $S_{h}^{2},\,h\in\mathcal{H}$. They also appear when one treats strata as domains and assigns upper bounds for variances of estimators of totals in domains. Such approach was considered e.g. in Choudhry et al. (2012), where apart of the upper bounds constraints $x_{h}\leq N_{h},\,h\in\mathcal{H}$, the additional constraints $\left(\tfrac{1}{x_{h}}-\tfrac{1}{N_{h}}\right)N_{h}^{2}S_{h}^{2}\leq R_{h},\,h\in\mathcal{H}$, where $R_{h},\,h\in\mathcal{H}$, are given constants, have been imposed. Obviously, the latter system of inequalities can be rewritten as lower bounds constraints of the form $x_{h}\geq m_{h}=\tfrac{N_{h}^{2}S_{h}^{2}}{R_{h}+N_{h}S_{h}^{2}},\,h\in\mathcal{H}$. The solution given in Choudhry et al. (2012) was obtained by the procedure based on the Newton-Raphson algorithm, a general-purpose root-finding numerical method. See also a related paper by Wright et al. (2007), where the optimum allocation problem under the constraint of the equal precision for estimation of the strata means was considered.
It is convenient to introduce the following definition for feasible solutions of Problem 1.1.
Definition 1.1.
Any vector $\mathbf{x}=(x_{h},\,h\in\mathcal{H})$ satisfying (1.4) and (1.5) will be called an allocation.
An allocation $\mathbf{x}=(x_{h},\,h\in\mathcal{H})$ is called a vertex one if and only if
[…]
where $\mathcal{L},\,\mathcal{U}\subseteq\mathcal{H}$ are such that $\mathcal{L}\cup\mathcal{U}=\mathcal{H}$ and $\mathcal{L}\cap\mathcal{U}=\emptyset$.
An allocation which is not a vertex one will be called a regular allocation.
The solution to Problem 1.1 will be called the optimum allocation.
Note that an optimum allocation may be of a vertex or of a regular form. The name vertex allocation refers to the fact that in this case $\mathbf{x}$ is a vertex of the hyper-rectangle ×$\times_{h\in\mathcal{H}}\,[m_{h},\,M_{h}]$. We note that Problem 1.1 becomes trivial if $n=\sum_{h\in\mathcal{H}}\,m_{h}$ or $n=\sum_{h\in\mathcal{H}}\,M_{h}$. In the former case, the solution is $\mathbf{x}^{*}=(m_{h},\,h\in\mathcal{H})$, and in the latter $\mathbf{x}^{*}=(M_{h},\,h\in\mathcal{H})$. These two are boundary cases of the vertex allocation. In real surveys with many strata, a vertex optimum allocation rather would not be expected. Nevertheless, for completeness we also consider such a case in Theorem 3.1, which describes the form of the optimum allocation vector. We also note that a regular optimum allocation $\mathbf{x}^{*}\in\times_{h\in\mathcal{H}}\,(m_{h},\,M_{h})$ if and only if it is the classical Tschuprow-Neyman allocation $\mathbf{x}^{*}=(A_{h}\,\tfrac{n}{\sum_{v\in\mathcal{H}}A_{v}},\,h\in\mathcal{H})$ (see Tschuprow, 1923, Neyman, 1934).
The rest of this paper is structured as follows. Section 2 presents motivations for this research as well as a brief review of the literature. In Section 3, we identify Problem 1.1 as a convex optimization problem and then use the Karush-Kuhn-Tucker conditions to establish necessary and sufficient conditions for a solution to optimization Problem 1.1. These conditions, called the optimality conditions, are presented in Theorem 3.1. In Section 4, based on these optimality conditions, we introduce a new algorithm, RNABOX, and prove that it solves Problem 1.1 (see Theorem 4.1). The name RNABOX refers to the fact that this algorithm generalizes the recursive Neyman algorithm, denoted here RNA. The RNA is a well-established allocation procedure, commonly used in everyday survey practice. It finds a solution to the allocation Problem 2.1 (see below), which is a relaxed version of Problem 1.1. As we shall see in Section 4.4, a naive modification of the recursive Neyman approach, which works fine for the one-sided upper bounds imposed on the sample strata sizes, does not give the correct solution in the case of two-sided bounds. A more subtle approach is needed.
Finally, let us note that the implementation of RNABOX algorithm is available through our R package stratallo (Wójciak, 2023b), which is published in CRAN repository (R Core Team, 2023).
2 Motivation and literature review
An abundant body of literature is devoted to the problem of optimum sample allocation, going back to classical solution of Tschuprow (1923) and Neyman (1934), dedicated to STSI sampling without taking inequality constraints (1.5) into account. In spite of this fact, a thorough analysis of the literature shows that Problem 1.1 has not been completely understood yet and it suffers from the lack of fully satisfactory algorithms.
Below, we briefly review existing methods for solving Problem 1.1, including also methods that provide integer-valued solutions.
2.1 Not-necessarily integer-valued allocation
An approximate solution to Problem 1.1 can be achieved through generic methods of non-linear programming (NLP) (see, e.g. the monograph Valliant et al., 2018, and references therein). These methods have been involved in the problem of optimum sample allocation since solving the allocation problem is equivalent to finding the extreme (namely, stationary points) of a certain objective function over a feasible set. Knowing the (approximate) extreme of the objective function, one can give the (approximate, yet typically sufficiently accurate) sizes of the samples allocated to individual strata.
In a similar yet different approach adopted e.g. in Münnich et al. (2012), Problem 1.1 is transformed into root-finding or fixed-point-finding problems (of some properly defined function) to which the solution is obtained by general-purpose algorithms like e.g. bisection or regula falsi.
Algorithms used in both these approaches would in principle have infinitely many steps, and are stopped by an arbitrary decision, typically related to the precision of the iterates. There are two main weaknesses associated with this way of operating: failure of the method to converge or slow convergence towards the optimal solution for some poor starting points. In other words, performance of these algorithms may strongly depend on an initial choice of a starting point, and such a choice is almost always somewhat hazardous. These and similar deficiencies are discussed in details in Appendix A.2 in the context of the allocation methods described in Münnich et al. (2012). Another drawback of the algorithms of this type is their sensitivity to finite precision arithmetic issues that can arise in case when the stopping criterion is not expressed directly in terms of the allocation vector iterates (which is often the case).
Contrary to that, in the recursive algorithms (we are concerned with), the optimal solution is always found by recursive search of feasible candidates for the optimum allocation among subsets of $\mathcal{H}$. Hence, they stop always at the exact solution and after finitely many iterations (not exceeding the number of strata + 1, as we will see for the case of RNABOX in the proof of Theorem 4.1). An important example of such an algorithm, is the recursive Neyman algorithm, RNA, dedicated for Problem 2.1, a relaxed version of Problem 1.1.
Problem 2.1.
Given a finite set $\mathcal{H}\neq\emptyset$ and numbers $A_{h}>0,\,M_{h},\,n>0$, such that $0<M_{h}\leq N_{h},\,h\in\mathcal{H}$ and $n\leq\sum_{h\in\mathcal{H}}\,M_{h}$,
$\underset{\mathbf{x}\,=\,(x_{h},\,h\,\in\,\mathcal{H})\,\in\,{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}}{\mathrm{minimize\leavevmode\nobreak\ \,}}$ $\quad\sum_{h\in\mathcal{H}}\tfrac{A_{h}^{2}}{x_{h}}$
$\mathrm{subject\leavevmode\nobreak\ to\quad\,\,\,}$ $\quad\sum_{h\in\mathcal{H}}x_{h}=n$
$\quad x_{h}\leq M_{h},\mathrlap{\qquad h\in\mathcal{H}.}$
Although RNA is popular among practitioners, a formal proof of the fact that it gives the optimal solution to Problem 2.1 has been given only recently in Wesołowski et al. (2022). For other recursive approaches to Problem 2.1, see also e.g. Stenger and Gabler (2005), Kadane (2005).
To the best of our knowledge, the only non-integer recursive optimum allocation algorithm described in the literature that is intended to solve Problem 1.1 is the noptcond procedure proposed by Gabler et al. (2012). In contrary to RNABOX, this method in particular performs strata sorting. Unfortunately, the allocation computed by noptcond may not yield the minimum of the objective function (1.3); see Appendix A.1 for more details.
2.2 Integer-valued allocation
Integer-valued algorithms dedicated to Problem 1.1 are proposed in Friedrich et al. (2015), Wright (2017, 2020). The multivariate version of the optimum sample allocation problem under box constraints in which $m_{h}=m,\,h\in\mathcal{H}$, for a given constant $m$, is considered in the paper of de Moura Brito et al. (2015). The proposed procedure that solves that problem, uses binary integer programming algorithm and can be applied to the univariate case. See also Brito et al. (2017) for the R-implementation of this approach.
Integer-valued allocation methods proposed in these papers are precise (not approximate) and theoretically sound. However, they are relatively slow, when compared with not-necessarily integer-valued algorithms. For instance, at least for one-sided constraints, the integer-valued algorithm capacity scaling of Friedrich et al. (2015) may be thousands of times slower than the RNA (see Wesołowski et al., 2022, Section 4). This seems to be a major drawback of these methods as the differences in variances of estimators based on integer-rounded non-integer optimum allocation and integer optimum allocation are negligible as explained in Section 6. The computational efficiency is of particular significance when the number of strata is large, see, e.g. application to the German census in Burgard and Münnich (2012), and it becomes even more pronounced in iterative solutions to stratification problems, when the number of iterations may count in millions (see, e.g. Lednicki and Wieczorkowski, 2003, Gunning and Horgan, 2004, Ahmad et al., 2008, Baillargeon and Rivest, 2011, Barcaroli, 2014).
Having all that said, the search for a new, universal, theoretically sound and computationally effective recursive algorithms of optimum sample allocation under two-sided constraints on the sample strata-sizes, is crucial both for theory and practice of survey sampling.
3 Optimality conditions
In this section we establish optimality conditions, that is, a general form of the solution to Problem 1.1. As it will be seen in Section 4, these optimality conditions are crucial for the construction of RNABOX algorithm.
Before we establish necessary and sufficient optimality conditions for a solution to optimization Problem 1.1, we first define a set function $s$, which considerably simplifies notation and calculations.
Definition 3.1.
Let $\mathcal{H},\,n,\,A_{h}>0,\,m_{h},\,M_{h},\,h\in\mathcal{H}$ be as in Problem 1.1 and let $\mathcal{L},\,\mathcal{U}\subseteq\mathcal{H}$ be such that $\mathcal{L}\cap\mathcal{U}=\emptyset,\,\mathcal{L}\cup\mathcal{U}\subsetneq\mathcal{H}$. The set function $s$ is defined as
$s(\mathcal{L},\,\mathcal{U})=\frac{n-\sum_{h\in\mathcal{L}}m_{h}-\sum_{h\in\mathcal{U}}M_{h}}{\sum_{h\in\mathcal{H}}A_{h}-\sum_{h\in\mathcal{L}\cup\mathcal{U}}A_{h}}.$ (3.1)
Below, we will introduce the $\mathbf{x}^{(\mathcal{L},\,\mathcal{U})}$ vector for disjoint $\mathcal{L},\,\mathcal{U}\subseteq\mathcal{H}$. It appears that the solution of the Problem 1.1 is necessarily of the form (3.2) with sets $\mathcal{L}$ and $\mathcal{U}$ defined implicitly through systems of equations/inequalities established in Theorem 3.1.
Definition 3.2.
Let $\mathcal{H},\,n,\,A_{h}>0,\,m_{h},\,M_{h},\,h\in\mathcal{H}$ be as in Problem 1.1, and let $\mathcal{L},\,\mathcal{U}\subseteq\mathcal{H}$ be such that $\mathcal{L}\cap\mathcal{U}=\emptyset$. We define the vector $\mathbf{x}^{(\mathcal{L},\,\mathcal{U})}=(x_{h}^{(\mathcal{L},\,\mathcal{U})},\,h\in\mathcal{H})$ as follows
[…]
The following Theorem 3.1 characterizes the form of the optimal solution to Problem 1.1 and therefore is one of the key results of this paper.
Theorem 3.1 (Optimality conditions).
The optimization Problem 1.1 has a unique optimal solution. Point $\mathbf{x}^{*}\in{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}$ is a solution to optimization Problem 1.1 if and only if $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$, with disjoint *$\mathcal{L}^{*},\,\mathcal{U}^{*}\subseteq\mathcal{H}$, such that one of the following two cases holds:
  1. CASE I:
    *$\mathcal{L}^{*}\cup\mathcal{U}^{*}\subsetneq\mathcal{H}$ and
    *=$\begin{gathered}\mathcal{L}^{*}=\left\{h\in\mathcal{H}:\,s(\mathcal{L}^{*},\,\mathcal{U}^{*})\leq\tfrac{m_{h}}{A_{h}}\right\},\\ \mathcal{U}^{*}=\left\{h\in\mathcal{H}:\,s(\mathcal{L}^{*},\,\mathcal{U}^{*})\geq\tfrac{M_{h}}{A_{h}}\right\}.\end{gathered}$ (3.3)
  2. CASE II:
    *$\mathcal{L}^{*}\cup\mathcal{U}^{*}=\mathcal{H}$ and
    $\max_{h\in\mathcal{U}^{*}}\tfrac{M_{h}}{A_{h}}\leq\min_{h\in\mathcal{L}^{*}}\tfrac{m_{h}}{A_{h}}\mathrlap{\qquad\text{ if\, }\mathcal{U}^{*}\neq\emptyset\text{ and }\mathcal{L}^{*}\neq\emptyset,},$ (3.4)
    $\sum_{h\in\mathcal{L}^{*}}m_{h}+\sum_{h\in\mathcal{U}^{*}}M_{h}=n.$ (3.5)
Remark 3.1.
The optimum allocation $\mathbf{x}^{*}$ is a regular one in CASE I and a vertex one in CASE II.
The proof of Theorem 3.1 is given in Appendix B. Note that Theorem 3.1 describes the general form of the optimum allocation up to specification of take-min and take-max strata sets $\mathcal{L}^{*}$ and $\mathcal{U}^{*}$. The question how to identify sets $\mathcal{L}^{*}$ and $\mathcal{U}^{*}$ that determine the optimal solution $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ is the subject of Section 4.
4 Recursive Neyman algorithm under box constraints
4.1 The RNABOX algorithm
In this section we introduce an algorithm solving Problem 1.1. In view of Theorem 3.1 its essential task is to split the set of all strata labels $\mathcal{H}$ into three subsets of take-min ($\mathcal{L}^{*}$), take-max ($\mathcal{U}^{*}$), and take-Neyman ((*$\mathcal{H}\setminus(\mathcal{L}^{*}\cup\mathcal{U}^{*})$). We call this new algorithm RNABOX since it generalizes existing algorithm RNA in the sense that RNABOX solves optimum allocation problem with simultaneous lower and upper bounds, while the RNA is dedicated for the problem with upper bounds only, i.e. for Problem 2.1. Moreover, RNABOX uses RNA in one of its interim steps. We first recall RNA algorithm and then present RNABOX. For more information on RNA, see Wesołowski et al. (2022, Section 2) or Särndal et al. (1992, Remark 12.7.1, p. 466).
Algorithm RNA
Input: $\mathcal{H},\,(A_{h})_{h\in\mathcal{H}},\,(M_{h})_{h\in\mathcal{H}},\,n$.
[…]
Step 2:Set $\mathcal{U}=\emptyset$.
Step 3:Determine $\widetilde{\mathcal{U}}=\left\{h\in\mathcal{H}\setminus\mathcal{U}:\,A_{h}\,s(\emptyset,\,\mathcal{U})\geq M_{h}\right\}$, where set function $s$ is defined in (3.1).
Step 4:If $\widetilde{\mathcal{U}}=\emptyset$, go to Step 5 . Otherwise, update $\mathcal{U}\leftarrow\mathcal{U}\cup\widetilde{\mathcal{U}}$ and go to Step 3 .
Step 5:Return $\mathbf{x}^{*}=(x^{*}_{h},\,h\in\mathcal{H})$ with $x^{*}_{h}=\begin{cases}M_{h},&h\in\mathcal{U}\\ A_{h}\,s(\emptyset,\,\mathcal{U}),&h\in\mathcal{H}\setminus\mathcal{U}.\end{cases}$
Algorithm RNABOX
Input: $\mathcal{H},\,(A_{h})_{h\in\mathcal{H}},\,(m_{h})_{h\in\mathcal{H}},\,(M_{h})_{h\in\mathcal{H}},\,n$.
[…]
Step 2:Set $\mathcal{L}=\emptyset$.
Step 3:Run RNA[$\mathcal{H},\,(A_{h})_{h\in\mathcal{H}},\,(M_{h})_{h\in\mathcal{H}},\,n$] to obtain $(x^{**}_{h},\,h\in\mathcal{H})$.
Let $\mathcal{U}=\{h\in\mathcal{H}:\,x^{**}_{h}=M_{h}\}$.
Step 4:Determine ~=$\widetilde{\mathcal{L}}=\left\{h\in\mathcal{H}\setminus\mathcal{U}:\,x^{**}_{h}\leq m_{h}\right\}$.
Step 5:If $\widetilde{\mathcal{L}}=\emptyset$ go to Step 6. Otherwise, update $n\leftarrow n-\sum_{h\in\widetilde{\mathcal{L}}}m_{h},\,\mathcal{H}\leftarrow\mathcal{H}\setminus\widetilde{\mathcal{L}}$, $\mathcal{L}\leftarrow\mathcal{L}\cup\widetilde{\mathcal{L}}$ and go to Step 3.
Step 6:Return $\mathbf{x}^{*}=(x^{*}_{h},\,h\in\mathcal{L}\cup\mathcal{H})$ with $x_{h}^{*}=\begin{cases}m_{h},&h\in\mathcal{L}\\ x^{**}_{h},&h\in\mathcal{H}.\end{cases}$
We note that in real life applications numbers $(A_{h})_{h\in\mathcal{H}}$ are typically unknown and therefore their estimates $(\hat{A}_{h})_{h\in\mathcal{H}}$ are used instead in the input of the algorithms.
Theorem 4.1 is the main theoretical result of this paper and its proof is given in Appendix C.
Theorem 4.1.
The RNABOX algorithm provides the optimal solution to Problem 1.1.
4.2 An example of performance of RNABOX
We demonstrate the operational behaviour of RNABOX algorithm for an artificial population with 10 strata and for total sample size $n=5110$, as shown in Table 4.1.
Table 4.1: An example of RNABOX performance for a population with 10 strata and total sample size $n=5110$. Columns $\mathcal{L}_{r}/\mathcal{U}_{r}$, $r=1,\ldots,6$, represent the content of sets $\mathcal{L},\,\mathcal{U}$ respectively, in the $r$-th iteration of the RNABOX (between Step 4 and Step 5): symbols $\blacksquare$  or $\square$  indicate that a given stratum label $h$ is in $\mathcal{L}_{r}$ or $\mathcal{U}_{r}$, respectively.
$h$
$A_{h}$ $m_{h}$ $M_{h}$ $\mathcal{L}_{1}/\mathcal{U}_{1}$ $\mathcal{L}_{2}/\mathcal{U}_{2}$ $\mathcal{L}_{3}/\mathcal{U}_{3}$ $\mathcal{L}_{4}/\mathcal{U}_{4}$ $\mathcal{L}_{5}/\mathcal{U}_{5}$ $\mathcal{L}_{6}/\mathcal{U}_{6}$ $\mathbf{x}^{*}$
1
2700 750 900 $\square$ $\square$ $\square$ $\square$ $750$
2
2000 450 500 $\blacksquare$ $\square$ $\square$ $\square$ $\square$ $450$
3
4200 250 300 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $261.08$
4
4400 350 400 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $350$
5
3200 150 200 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $198.92$
6
6000 550 600 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $\square$ $550$
7
8400 650 700 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $650$
8
1900 50 100 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $100$
9
5400 850 900 $\blacksquare$ $\blacksquare$ $\square$ $\square$ $\square$ $850$
10
2000 950 1000 $\square$ $\square$ $\square$ $\square$ $\square$ $950$
SUM 5000 5600 0/8 1/7 3/6 4/3 5/3 7/1 $5110$
For this example, RNABOX stops after $6$ iterations with take-min strata set *=$\mathcal{L}^{*}=\{1,2,4,6,7,9,10\}$, take-max strata set $\mathcal{U}^{*}=\{8\}$ and take-Neyman strata set (*$\mathcal{H}\setminus(\mathcal{L}^{*}\cup\mathcal{U}^{*})=\{3,5\}$ (see column $\mathcal{L}_{6}/\mathcal{U}_{6}$). The optimum allocation is a regular one and it is given in column $\mathbf{x}^{*}$ of Table 4.1. The corresponding value of the objective function (1.3) is $441591.5$. The details of interim allocations of strata to sets $\mathcal{L},\,\mathcal{U}$ at each of $6$ iterations of the algorithm are given in columns $\mathcal{L}_{1}/\mathcal{U}_{1}$ $\mathcal{L}_{6}/\mathcal{U}_{6}$.
Results summarized in Table 4.1 are presented graphically in Fig. 4.1.
Figure 4.1: Assignments of strata into set $\mathcal{L}$ (take-min) and set $\mathcal{U}$ (take-max) in RNABOX algorithm for an example of population as given in Table 4.1 and total sample size $n=5110$. The $\sigma(\mathcal{H})$ axis corresponds to strata assigned to set $\mathcal{U}$, while $\tau(\mathcal{H})$ is for $\mathcal{L}$. Squares represent assignments of strata to $\mathcal{L}$ ($\square$) or $\mathcal{U}$ ($\blacksquare$) such that the coordinate corresponding to a given square is the value of the last element (following the order of strata, $\sigma$ or $\tau$, associated to the respective axis) in the set.
The $\sigma(\mathcal{H})$ axis corresponds to strata assigned to set $\mathcal{U}$, while $\tau(\mathcal{H})$ is for $\mathcal{L}$. For the sake of clear presentation, the values on $\sigma(\mathcal{H})$ axis are ordered non-increasingly in $\tfrac{A_{h}}{M_{h}}$, $h\in\mathcal{H}$, which gives $\tfrac{A_{8}}{M_{8}}\geq\tfrac{A_{5}}{M_{5}}\geq\tfrac{A_{3}}{M_{3}}\geq\ldots\geq\tfrac{A_{10}}{M_{10}}$, while for $\tau(\mathcal{H})$ axis the strata are ordered non-decreasingly in $\tfrac{A_{h}}{m_{h}}$, $h\in\mathcal{H}$, which gives $\tfrac{A_{10}}{m_{10}}\leq\tfrac{A_{1}}{m_{1}}\leq\tfrac{A_{2}}{m_{2}}\leq\ldots\leq\tfrac{A_{8}}{m_{8}}$. Squares in Fig. 4.1 represent assignments of strata into sets $\mathcal{L}$ or $\mathcal{U}$, such that the corresponding coordinate is the value of the last element (follow the respective axis) in the set. For example, $\blacksquare$  at coordinate $6$ on $\sigma(\mathcal{H})$ axis means that in this iteration of the algorithm (iteration no. $3$) we have $\mathcal{U}=\{8,5,3,7,4,6\}$, while $\square$  at coordinate $6$ on $\tau(\mathcal{H})$ axis means that in this iteration (no. $5$) we get =$\mathcal{L}=\{10,1,2,9,6\}$. Directions indicated by the arrows correspond to the order in which the sets $\mathcal{L}$ and $\mathcal{U}$ are built as the algorithm iterates. Numbers above the squares are the values of set function $s(\mathcal{L},\,\mathcal{U})$ for $\mathcal{L}$ and $\mathcal{U}$ with elements indicated by the coordinates of a given square. For example, for $\blacksquare$  at coordinates (8, 7)87(8,\,7)( 8 , 7 ), we have $s(\{10,1,2,9,6,4,7\},\,\{8\})=0.0622$.
We would like to point out that regardless of the population and other allocation parameters chosen, the graph illustrating the operations of RNABOX algorithm will always have a shape similar to that of the right half of a Christmas tree with the top cut off when $\mathcal{U}_{r^{*}}\neq\emptyset$. This property results from that fact that $\mathcal{L}_{r}\subsetneq\mathcal{L}_{r+1}$(see (C.1)) and $\mathcal{U}_{r}\supseteq\mathcal{U}_{r+1}$ (see (C.16)) for $r=1,\ldots,r^{*}-1,\,r^{*}\geq 2$. For the population given in Table 4.1, we clearly see that $\mathcal{L}_{1}=\emptyset\subsetneq\mathcal{L}_{2}=\{10\}\subsetneq\mathcal{L}_{3}=\{10,1,2\}\subsetneq\mathcal{L}_{4}=\{10,1,2,9\}\subsetneq\mathcal{L}_{5}=\{10,1,2,9,6\}\subsetneq\mathcal{L}_{6}=\mathcal{L}^{*}=\{10,1,2,9,6,4,7\}$ and $\mathcal{U}_{1}=\{8,5,3,7,4,6,9,2\}\supset\mathcal{U}_{2}=\{8,5,3,7,4,6,9\}\supset\mathcal{U}_{3}=\{8,5,3,7,4,6\}\supset\mathcal{U}_{4}=\{8,5,3\}=\mathcal{U}_{5}\supset\mathcal{U}_{6}=\mathcal{U}^{*}=\{8\}$. Moreover, regardless of the population and other allocation parameters chosen, subsequent values of set function $s$, which are placed at the ends of branches of the Christmas tree, above $\blacksquare$, form a non-increasing sequence while moving upwards. This fact follows directly from Lemma C.3. For the example allocation illustrated in Fig. 4.1, we have $s(\mathcal{L}_{1},\,\mathcal{U}_{1})=0.3>s(\mathcal{L}_{2},\,\mathcal{U}_{2})=0.204>s(\mathcal{L}_{3},\,\mathcal{U}_{3})=0.122>s(\mathcal{L}_{4},\,\mathcal{U}_{4})=0.0803>s(\mathcal{L}_{5},\,\mathcal{U}_{5})=0.075>s(\mathcal{L}_{6},\,\mathcal{U}_{6})=0.0622$ (note that $\mathcal{L}_{1}=\emptyset$). The same property appears for values of $s$ related to the trunk of the Christmas tree, placed above $\square$. In this case, this property is due to (C.2), (C.15) and (C.1). For allocation in Fig. 4.1, we have $s(\mathcal{L}_{1},\,\emptyset)=0.127>s(\mathcal{L}_{2},\,\emptyset)=0.109>s(\mathcal{L}_{3},\,\emptyset)=0.0884>s(\mathcal{L}_{4},\,\emptyset)=0.0751>s(\mathcal{L}_{5},\,\emptyset)=0.0706>s(\mathcal{L}_{6},\,\emptyset)=0.0602$.
4.3 Possible modifications and improvements
4.3.1 Alternatives for RNA in Step 2
The RNABOX algorithm uses RNA in its Step 3. However, it is not hard to see that any algorithm dedicated to Problem 2.1 (like for instance SGA by Stenger and Gabler, 2005 or COMA by Wesołowski et al., 2022) could be used instead. We chose RNA as it allows to keep RNABOX free of any strata sorting.
4.3.2 A twin version of RNABOX
Let us observe that the order in which $\mathcal{L}$ and $\mathcal{U}$ are computed in the algorithm could be interchanged. Such a change, implies that the RNA used in Step 3 of the RNABOX, should be replaced by its twin version, the LRNA, that solves optimum allocation problem under one-sided lower bounds. The LRNA is described in details in Wójciak (2023a).
Algorithm LRNA
Input: $\mathcal{H},\,(A_{h})_{h\in\mathcal{H}},\,(m_{h})_{h\in\mathcal{H}},\,n$.
Step 1:$A_{h}>0,\,m_{h}>0,\,h\in\mathcal{H}$, $n\geq\sum_{h\in\mathcal{H}}\,m_{h}$.
Step 2:Let $\mathcal{L}=\emptyset$.
Step 3:Determine ~=$\widetilde{\mathcal{L}}=\left\{h\in\mathcal{H}\setminus\mathcal{L}:\,A_{h}\,s(\mathcal{L},\,\emptyset)\leq m_{h}\right\}$, where set function $s$ is defined in (3.1).
Step 4: If $\widetilde{\mathcal{L}}=\emptyset$, go to Step 5 . Otherwise, update $\mathcal{L}\leftarrow\mathcal{L}\cup\widetilde{\mathcal{L}}$ and go to Step 3 .
Step 5:Return $\mathbf{x}^{*}=(x^{*}_{h},\,h\in\mathcal{H})$ with $x^{*}_{h}=\begin{cases}m_{h},&h\in\mathcal{L}\\ A_{h}\,s(\mathcal{L},\,\emptyset),&h\in\mathcal{H}\setminus\mathcal{L}.\end{cases}$
Taking into account the observation above, Step 3 and Step 4 of RNABOX would read:
Step 2:  Run LRNA[$\mathcal{H},\,(A_{h})_{h\in\mathcal{H}},\,(m_{h})_{h\in\mathcal{H}},\,n$] to obtain $(x^{**}_{h},\,h\in\mathcal{H})$.
Let =$\mathcal{L}=\{h\in\mathcal{H}:\,x^{**}_{h}=m_{h}\}$.
Step 3:  Determine $\widetilde{\mathcal{U}}=\left\{h\in\mathcal{H}\setminus\mathcal{L}:\,x^{**}_{h}\geq M_{h}\right\}$.
The remaining steps should be adjusted accordingly.
4.3.3 Using prior information in RNA at Step 3
In view of Lemma C.2, using the notation introduced in Appendix C.1, in Step 3 of RNABOX, for $r^{*}\geq 2$ we have
$*.\mathcal{U}_{r}=\{h\in\mathcal{H}\setminus\mathcal{L}_{r}:\,x^{**}_{h}=M_{h}\}\subseteq\mathcal{U}_{r-1},\mathrlap{\qquad r=2,\ldots,r^{*}.}$
This suggests that the domain of discourse for $\mathcal{U}_{r}$ could be shrunk from $\mathcal{H}\setminus\mathcal{L}_{r}$ to $\mathcal{U}_{r-1}\subseteq\mathcal{H}\setminus\mathcal{L}_{r}$, i.e.
$*.\mathcal{U}_{r}=\{h\in\mathcal{U}_{r-1}:\,x^{**}_{h}=M_{h}\},\mathrlap{\qquad r=2,\ldots,r^{*}.}$ (4.1)
Given above observation and the fact that from the implementation point of view set $\mathcal{U}_{r}$ is determined internally by RNA, it is tempting to consider modification of RNA such that it makes use of the domain of discourse $\mathcal{U}_{r-1}$ for set $\mathcal{U}_{r}$. This domain could be specified as an additional input parameter, say $\mathcal{J}\subseteq\mathcal{H}$, and then Step 3 of RNA algorithm would read:
Step 2:  Determine $\widetilde{\mathcal{U}}=\left\{h\in\mathcal{J}\setminus\mathcal{U}:\,A_{h}\,s(\emptyset,\,\mathcal{U})\geq M_{h}\right\}$.
From RNABOX perspective, this new input parameter of RNA should be set to $\mathcal{J}=\mathcal{H}$ for the first iteration, and then $\mathcal{J}=\mathcal{U}_{r-1}$ for subsequent iterations $r=2,\ldots,r^{*}\geq 2$ (if any).
4.4 On a naive extension of the one-sided RNA
By the analogy to the one-sided constraint case, one would expect that the optimum allocation problem under two-sided constraints (1.5) is of the form (3.2) and hence it could be solved with the naive modification of RNA as defined below.
Algorithm Naive modification of RNA
Input: $\mathcal{H},\,(A_{h})_{h\in\mathcal{H}},\,(m_{h})_{h\in\mathcal{H}},\,(M_{h})_{h\in\mathcal{H}},\,n$.
[…]
Step 2:Set $\mathcal{L}=\emptyset,\,\mathcal{U}=\emptyset$.
Step 3:Determine $\widetilde{\mathcal{U}}=\{h\in\mathcal{H}\setminus(\mathcal{L}\cup\mathcal{U}):\,A_{h}\,s(\mathcal{L},\,\mathcal{U})\geq M_{h}\}$, where set function $s$ is defined in (3.1).
Step 4:Determine ~=$\widetilde{\mathcal{L}}=\{h\in\mathcal{H}\setminus(\mathcal{L}\cup\mathcal{U}):\,A_{h}\,s(\mathcal{L},\,\mathcal{U})\leq m_{h}\}$.
Step 5:If ~$\widetilde{\mathcal{L}}\cup\widetilde{\mathcal{U}}=\emptyset$ go to Step 6. Otherwise, update ~$\mathcal{L}\leftarrow\mathcal{L}\cup\widetilde{\mathcal{L}},\,\mathcal{U}\leftarrow\mathcal{U}\cup\widetilde{\mathcal{U}}$ and go to Step 3.
Step 6:Return $\mathbf{x}^{*}=(x^{*}_{h},\,h\in\mathcal{H})$ with $x^{*}_{h}=\begin{cases}m_{h},&\leavevmode\nobreak\ h\in\mathcal{L}\\ M_{h},&\leavevmode\nobreak\ h\in\mathcal{U}\\ A_{h}\,s(\mathcal{L},\,\mathcal{U}),&\leavevmode\nobreak\ h\in\mathcal{H}\setminus(\mathcal{L}\cup\mathcal{U}).\end{cases}$
This is however not true and the basic reason behind is that sequence $\left(s(\mathcal{L}_{r},\,\mathcal{U}_{r})\right)_{r=1}^{r^{*}}$ (where $r$ denotes iteration index, and $\mathcal{L}_{r}$ and $\mathcal{U}_{r}$ are taken after Step 4 and before Step 5), typically fluctuates. Thus, it may happen that (C.13) - (C.15) are not met for $\mathcal{U}_{r},\,\widetilde{\mathcal{L}}_{r}$ defined by the above naive algorithm for some $r=1,\ldots,r^{*}\geq 1$. Consequently, optimality conditions stated in Theorem 3.1 may not hold. To illustrate this fact, consider an example of population as given in Table 4.2.
[…]
$h$
[…]
1
420 24 420 $30$ $54.44$
2
352 15 88 $88$blacksquare $45.63$
3
2689 1344 2689 $1344$square $1344$square
4
308 8 308 $22$ $39.93$
5
130 3 5 $5$blacksquare $5$blacksquare
SUM
1394 3510 $4891$/2/2 $4891$/1/3
For this population, a naive version of the recursive Neyman stops in the second iteration and it gives allocation $\mathbf{x}$ with =$\mathcal{L}=\{3\}$ and $\mathcal{U}=\{2,5\}$, while the optimum choice is $\mathbf{x}^{*}$ with *=$\mathcal{L}^{*}=\{3\}$ and $\mathcal{U}^{*}=\{5\}$. For non-optimum allocation $\mathbf{x}$, we have $s(\{3\},\,\{2,5\})=0.0714\ngeq 0.25=\tfrac{A_{2}}{M_{2}}$, i.e. $\mathcal{U}$ does not follow (3.3). Note that the allocation $\mathbf{x}$ is feasible in this case (it may not be in general), yet is not a minimizer of the objective function (1.3) as $f(\mathbf{x})=20360-B$, and $f(\mathbf{x}^{*})=17091-B$, where $B$ is some constant. Thus, it appears that in the case of two-sided constraints, one needs a more subtle modification of RNA procedure that lies in a proper update of the pair $(\mathcal{L}_{r},\,\mathcal{U}_{r})$ in each iteration $r=1,\ldots,r^{*}\geq 1$ of the algorithm, such as e.g. in RNABOX.
5 Numerical results
In simulations, using R Statistical Software (R Core Team, 2023) and microbenchmark R package (Mersmann, 2021), we compared the computational efficiency of RNABOX algorithm with the efficiency of the fixed-point iteration algorithm (FPIA) of Münnich et al. (2012). The latter one is known to be an efficient algorithm dedicated to Problem 1.1 and therefore we used it as a benchmark. The comparison was not intended to verify theoretical computational complexity, but was rather concerned with quantitative results regarding computational efficiency for the specific implementations of both algorithms.
To compare the performance of the algorithms we used the STSI sampling for several artificially created populations. Here, we chose to report on simulation for two such populations with 691 and 703 strata, results of which are representative for the remaining ones. These two populations were constructed by iteratively binding $K$ sets of numbers, where $K$ equals 100 (for the first population) and 200 (for the second population). Each set, labelled by $i=1,\dots,K$, contains 10000 random numbers generated independently from log-normal distribution with parameters $\mu=0$ and $\sigma=\log(1+i)$. For every set $i=1,\dots,K$, strata boundaries were determined by the geometric stratification method of Gunning and Horgan (2004) with parameter 10 being the number of strata and targeted coefficient of variation equal to 0.05. This stratification method is implemented in the R package stratification, developed by Rivest and Baillargeon (2022) and described in Baillargeon and Rivest (2011). For more details, see the R code with the experiments, which is placed in our GitHub repository (see Wieczorkowski et al., 2023).
Results of these simulations are illustrated in Fig. 5.1.
Figure 5.1: Running times of FPIA and RNABOX for two artificial populations. Top graphs show the empirical median of execution times (calculated from 100 repetitions) for different total sample sizes. Numbers in brackets are the numbers of iterations of a given algorithm. In the case of RNABOX, it is a vector with number of iterations of the RNA (see Step 3 of RNABOX) for each outer iteration of RNABOX. Thus, the length of this vector is equal to the number of outer iterations of RNABOX. Counts of take-min, take-Neyman, and take-max strata are shown on bottom graphs.
From Fig. 5.1 we see that, while for majority of the cases the FPIA is slightly faster than RNABOX, the running times of both of these algorithms are generally comparable. The gain in the execution time of the FPIA results from the fact that it typically runs through a smaller number of sets $\mathcal{L},\,\mathcal{U}\subset\mathcal{H}$ (according to (A.6)), than RNABOX in order to find the optimal $\mathcal{L}^{*}$ and $\mathcal{U}^{*}$. Although this approach usually gives correct results (as in the simulations reported in this section), it may happen that the FPIA misses the optimal sets *$\mathcal{L}^{*},\,\mathcal{U}^{*}\subsetneq\mathcal{H}$. In such a case, FPIA does not converge, as explained in Appendix A.2. Nevertheless, we emphasize that this situation rarely happens in practice.
6 Concluding comments
In this paper we considered Problem 1.1 of optimum sample allocation under box constraints. The main result of this work is the mathematically precise formulation of necessary and sufficient conditions for the solution to Problem 1.1, given in Theorem 3.1, as well as the development of the new recursive algorithm, termed RNABOX, that solves Problem 1.1. The optimality conditions are fundamental to analysis of the optimization problem. They constitute trustworthy underlay for development of effective algorithms and can be used as a baseline for any future search of new algorithms solving Problem 1.1. Essential properties of RNABOX algorithm, that distinguish it from other existing algorithms and approaches to the Problem 1.1, are:
  1. 1.
    Universality: RNABOX provides optimal solution to every instance of feasible Problem 1.1 (including the case of a vertex optimum allocation).
  2. 2.
    No initialization issues: RNABOX does not require any initializations, pre-tests or whatsoever that could have an impact on the final results of the algorithm. This, in turn, takes places e.g. in case of NLP methods.
  3. 3.
    No sorting: RNABOX does not perform any ordering of strata.
  4. 4.
    Computational efficiency: RNABOX running time is comparable to that of FPIA (which is probably the fastest previously known optimum allocation algorithm for the problem considered).
  5. 5.
    Directness: RNABOX computes important quantities (including RNA internals) via formulas that are expressed directly in terms of the allocation vector $\mathbf{x}^{(\mathcal{L},\,\mathcal{U})}$ (see Definition 3.2). This reduces the risk of finite precision arithmetic issues, comparing to the algorithms that base their key operations on some interim variables on which the optimum allocation depends, as is the case of e.g. the NLP-based method.
  6. 6.
    Recursive nature: RNABOX repeatedly applies allocation Step 3 and Step 4 to step-wise reduced set of strata, i.e. "smaller" versions of the same problem. This translates to clarity of the routines and a natural way of thinking about the allocation problem.
  7. 7.
    Generalization: RNABOX, from the perspective of its construction, is a generalization of the popular RNA algorithm that solves Problem 2.1 of optimum sample allocation under one-sided bounds on sample strata sizes.
Finally, we would like to note that Problem 1.1 considered in this paper is not an integer-valued allocation problem, while the sample sizes in strata should be of course natural numbers. On the other hand, the integer-valued optimum allocation algorithms are relatively slow and hence might be inefficient in some applications, as already noted in Section 2. If the speed of an algorithm is of concern and non-necessarily integer-valued allocation algorithm is chosen (e.g. RNABOX), the natural remedy is to round the non-integer optimum allocation provided by that algorithm. Altogether, such procedure is still much faster than integer-valued allocation algorithms. However, a simple rounding of the non-integer solution does not, in general, yield the minimum of the objective function, and may even lead to an infeasible solution, as noted in Friedrich et al. (2015, Section 1, p. 3). Since infeasibility can in fact arise only from violating constraint (1.4), it can be easily avoided by using a rounding method of Cont and Heidari (2014) that preserves the integer sum of positive numbers. Moreover, all numerical experiments that we carried out, show that the values of the objective function obtained for non-integer optimum allocation before and after rounding and for the integer optimum allocation are practically indistinguishable. See Table 6.1 for the exact numbers.
Table 6.1: For populations used in simulations in Section 5, $V_{int}$ and $V$ denote values of variance (1.1) computed for integer optimum allocations and non-integer optimum allocations, respectively. Variances $V_{round}$ are computed for rounded non-integer optimum allocations (with the rounding method of Cont and Heidari (2014)).
691 strata ($691\text{ strata }\leavevmode\nobreak\ (N=990403)$ 703 strata ($703\text{ strata }\leavevmode\nobreak\ (N=991226)$
fraction f $n=\text{ f }\cdot N$ $V/V_{int}$ $V_{round}/V_{int}$ $n=\text{ f }\cdot N$ $V/V_{int}$ $V_{round}/V_{int}$
0.1 99040 0.999997 1.000000 99123 0.999997 1.00000
0.2 198081 0.999999 1.000000 198245 0.999999 1.00000
0.3 297121 0.999999 1.000000 297368 0.999999 1.00000
0.4 396161 0.999999 1.000000 396490 0.999999 1.00000
0.5 495202 0.999999 1.000000 495613 0.999999 1.00000
0.6 594242 0.999999 1.000000 594736 0.999999 1.00000
0.7 693282 1.000000 1.000000 693858 0.999998 1.00000
0.8 792322 1.000000 1.000000 792981 0.999995 1.00000
0.9 891363 1.000000 1.000000 892103 0.999759 1.00000
The above observations suggest that fast, not-necessarily integer-valued allocation algorithms, with properly rounded results, may be a good and reasonable alternative to slower integer algorithms when speed of an algorithm is crucial.
Appendix A Appendix: On the existing allocation algorithms by GGM and MSW
A.1 Necessary vs. sufficient optimality conditions and the noptcond function of GGM
Problem 1.1 has been considered in Gabler et al. (2012) (GGM in the sequel). Theorem 1 in that paper announces that there exist disjoint sets $\mathcal{L},\,\mathcal{U}\subset\mathcal{H}$ with $\mathcal{L}^{c}=\mathcal{H}\setminus\mathcal{L},\,\mathcal{U}^{c}=\mathcal{H}\setminus\mathcal{U}$, such that
$\max_{h\in\mathcal{L}}\tfrac{A_{h}}{m_{h}}<\min_{h\in\mathcal{L}^{c}}\tfrac{A_{h}}{m_{h}}\qquad\mbox{and}\qquad\max_{h\in\mathcal{U}^{c}}\tfrac{A_{h}}{M_{h}}<\min_{h\in\mathcal{U}}\tfrac{A_{h}}{M_{h}},$ (A.1)
and, if the allocation $\mathbf{x}=(x_{h},\,h\in\mathcal{H})$ is of the form (3.2) with $\mathcal{L}$ and $\mathcal{U}$ as above, then $\mathbf{x}$ is an optimum allocation. An explanation on how to construct sets $\mathcal{L}$ and $\mathcal{U}$ is essentially given in the first part of the proof of Theorem 1, p. 154-155, or it can be read out from the noptcond function. The noptcond is a function in R language (see R Core Team, 2023) that was defined in Sec. 3 of GGM and it aims to find an optimum sample allocation based on results of Theorem 1, as the authors pointed out. However, as it will be explained here, (3.2) and (A.1) together, are in fact necessary but not sufficient conditions for optimality of the allocation $\mathbf{x}$. In a consequence, the noptcond function may not give an optimum allocation.
By Theorem 3.1, CASE I, a regular allocation $\mathbf{x}$ is an optimum allocation if and only if it is of the form (3.2) with disjoint sets $\mathcal{L},\,\mathcal{U}\subsetneq\mathcal{H}$, $\mathcal{L}\cup\mathcal{U}\subsetneq\mathcal{H}$, $\mathcal{L}^{c}=\mathcal{H}\setminus\mathcal{L},\,\mathcal{U}^{c}=\mathcal{H}\setminus\mathcal{U}$ such that the following inequalities hold true:
$\max_{h\in\mathcal{L}}\tfrac{A_{h}}{m_{h}}\leq\tfrac{1}{s(\mathcal{L},\,\mathcal{U})}<\min_{h\in\mathcal{L}^{c}}\tfrac{A_{h}}{m_{h}},$ (A.2)
$\max_{h\in\mathcal{U}^{c}}\tfrac{A_{h}}{M_{h}}<\tfrac{1}{s(\mathcal{L},\,\mathcal{U})}\leq\min_{h\in\mathcal{U}}\tfrac{A_{h}}{M_{h}}.$ (A.3)
Clearly, (A.2) implies the first inequality in (A.1), and (A.3) implies the second one. Converse implications do not necessary hold, and therefore sufficiency of condition (3.2) with (A.1) for Problem 1.1 is not guaranteed. In others words, (3.2) and (A.1) alone, lead to a feasible solution (i.e. the one that does not violate any of the constraints (1.4) - (1.5)), which, at the same time, might not be a minimizer of the objective function (1.3).
The proof of Theorem 1 in GGM gives an explicit, but somewhat informal algorithm that finds sets $\mathcal{L}$ and $\mathcal{U}$ which define allocation (3.2). Sets $\mathcal{L}$ and $\mathcal{U}$ determined by this algorithm meet inequalities (4) and (5) in GGM, which are necessary and sufficient conditions for an optimal (regular) solution. This fact was noted by the authors, but unfortunately (4) and (5) are not given in the formulation of Theorem 1. Presumably, this was due to authors’ statement that (4) and (5) follow from (3.2) and (A.1): "From the definition of L1, L2 in Eq. 3 we have …", while in fact to conclude (4) and (5) the authors additionally rely on the algorithm constructed at the beginning of the proof. Note that conditions (4) and (5) are the same as those given in Theorem 3.1 for a regular solution (CASE I). Moreover, it is a matter of simple observation to see that the allocation found by the algorithm described in the proof of Theorem 1 in GGM, meets also optimality conditions (3.4) and (3.5) established in Theorem 3.1 for vertex solutions (CASE II). Hence, the allocation computed by the algorithm embedded in the proof of the GGM’s Theorem 1 is indeed an optimal one, though the formal statement of Theorem 1 is not correct.
Unfortunately, the noptcond function does not fully follow the algorithm from the proof of Theorem 1 of GGM. Consequently, the optimality of a solution computed by noptcond is not guaranteed. This can be illustrated by a simple numerical Example A.1, which follows Wójciak (2019, Example 3.9).
Example A.1.
Consider the allocation for an example population as given in Table A.1.
Table A.1: Two allocations for an example population with two strata: non-optimum $\mathbf{x}^{noptcond}$ with =$\mathcal{L}=\{1\},\,\mathcal{U}=\emptyset$, and optimum $\mathbf{x}^{*}$ with *$\mathcal{L}^{*}=\emptyset,\,\mathcal{U}^{*}=\{1\}$. Set function $s$ is defined in (3.1).
$h$ $A_{h}$ $m_{h}$ $M_{h}$ $\tfrac{A_{h}}{m_{h}}$ $\tfrac{A_{h}}{M_{h}}$ $s^{-1}(\{1\},\emptyset)$ $s^{-1}(\emptyset,\{1\})$ $\mathbf{x}^{noptcond}$ $\mathbf{x}^{*}$
1 2000 30 50 $66.67$ $40$ 23.08 27.27 $30$ $50$
2 3000 40 200 $75$ $15$ $130$ $110$
[…]
$\max_{h\in\mathcal{L}}\tfrac{A_{h}}{m_{h}}\leq\tfrac{A_{v}}{x_{v}}\leq\min_{h\in\mathcal{U}}\tfrac{A_{h}}{M_{h}},\mathrlap{\qquad v\in\mathcal{H}\setminus(\mathcal{L}\cup\mathcal{U}).}$ (A.4)
We finally note that the table given in Section 4 of GGM on p. 160, that illustrates how noptcond operates, is incorrect as some of the numbers given do not correspond to what noptcond internally computes, e.g. for each row of this table, the allocation should sum up to $n=20$, but it does not.
A.2 Fixed-point iteration of MSW
Theorem 1 of Münnich et al. (2012) (referred to by MSW in the sequel) provides another solution to Problem 1.1 that is based on results partially similar to those stated in Theorem 3.1. Specifically, it states that the allocation vector $\mathbf{x}=(x_{h},\,h\in\mathcal{H})$ can be expressed as a function $\mathbf{x}:\,{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}$ of $\lambda$, defined as follows
$x_{h}(\lambda)=\begin{cases}M_{h},&\qquad{h\in J^{\lambda}_{M}:=\left\{h\in\mathcal{H}:\,\lambda\leq\tfrac{A_{h}^{2}}{M_{h}^{2}}\right\}}\\ \tfrac{A_{h}}{\sqrt{\lambda}},&\qquad{h\in J^{\lambda}:=\left\{h\in\mathcal{H}:\,\tfrac{A_{h}^{2}}{M_{h}^{2}}<\lambda<\tfrac{A_{h}^{2}}{m_{h}^{2}}\right\}}\\ m_{h},&\qquad{h\in J^{\lambda}_{m}:=\left\{h\in\mathcal{H}:\,\lambda\geq\tfrac{A_{h}^{2}}{m_{h}^{2}}\right\},}\end{cases}$
and the optimum allocation $\mathbf{x}^{*}$ is obtained for $\lambda=\lambda^{*}$, where $\lambda^{*}$ is the solution of the equation:
${\tilde{g}}(\lambda):=\sum_{h\in\mathcal{H}}\,x_{h}(\lambda)-n=0.$ (A.5)
Here, we note that the optimum allocation $\mathbf{x}^{*}$ obtained is of the form $\mathbf{x}^{(\mathcal{L},\,\mathcal{U})}$, as given in (3.2), with $\mathcal{L}=J^{\lambda^{*}}_{m}$ and $\mathcal{U}=J^{\lambda^{*}}_{M}$. Since $\tilde{g}$ is continuous but not differentiable, equation (A.5) has to be solved by a root-finding methods which only require continuity. In MSW, the following methods were proposed: bisection, secan and regula falsi in few different versions. The authors reported that despite the fact that some theoretical requirements for these methods are not satisfied by the function $\tilde{g}$ (e.g. it is not convex while regula falsi applies to convex functions), typically the numerical results are satisfactory. However this is only true when the values of the initial parameters are from the proper range, which is not known a priori.
As these algorithms might be relatively slow (see Table 1 in MSW), the authors of MSW proposed an alternative method. It is based on the observation that (A.5) is equivalent to
$\lambda=\phi(\lambda):=\tfrac{1}{s^{2}(J_{m}^{\lambda},\,J^{\lambda}_{M})},$
where set function $s$ is as in (3.1) (see Appendix A, Lemma 2 in MSW). Note that $\phi$ is well-defined only if $J_{m}^{\lambda}\cup J_{M}^{\lambda}\subsetneq\mathcal{H}$. The authors note that the optimal $\lambda^{*}$ is a fixed-point of the function ϕ:++:italic-$\phi:{\mathbb{R}}_{+}\to{\mathbb{R}}_{+}$. This clever observation was translated to an efficient optimum allocation fixed-point iteration algorithm, FPIA, that was defined on page 442 in MSW. Having acknowledged its computational efficiency as well as the fact that typically it gives the correct optimum allocation, the following two minor issues related to FPIA can be noted:
  • The FPIA is adequate only for allocation problems for which an optimum allocation is of a regular type (according to Definition 1.1), as for vertex allocation $J_{m}^{\lambda^{*}}\cup J_{M}^{\lambda^{*}}=\mathcal{H}$ and therefore $\phi$ is not well-defined for such $\lambda^{*}$.
  • The FPIA strongly depends on the choice of initial value $\lambda_{0}$ of the parameter $\lambda$. If $\lambda_{0}$ is not chosen from the proper range (not known a priori), two distinct undesirable scenarios can occur: the FPIA may get blocked, or it may not converge.
To outline the blocking scenario which may happen even in case of a regular optimum allocation, let
$I:=\left\{\lambda\in\left[\min_{h\in\mathcal{H}}\tfrac{A^{2}_{h}}{m^{2}_{h}},\,\max_{h\in\mathcal{H}}\tfrac{A^{2}_{h}}{M^{2}_{h}}\right]:\,s(J_{m}^{\lambda},\,J^{\lambda}_{M})=0\right\},$
and note that $I\neq\emptyset$ is possible. If FPIA encounters $\lambda_{k}\in I\neq\emptyset$ at some $k=0,1,\ldots$, then the crucial step of FPIA, i.e.:
$\lambda_{k+1}:=\tfrac{1}{s^{2}\left(J_{m}^{\lambda_{k}},\,J^{\lambda_{k}}_{M}\right)},$ (A.6)
yields $\lambda_{k+1}=\tfrac{1}{0}$, which is undefined. Numerical Example A.2 illustrates this scenario.
Example A.2.
[…]
Table A.2: Details of FPIA performance for a population with four strata. Here, $\lambda_{0}=\tfrac{1}{s^{2}(\emptyset,\,\emptyset)},\,\lambda_{1}=\tfrac{1}{s^{2}(J_{m}^{\lambda_{0}},\,J_{M}^{\lambda_{0}})},\,\lambda^{*}=\tfrac{1}{s^{2}(J_{m}^{\lambda^{*}},\,J_{M}^{\lambda^{*}})}$. Squares in cells indicate assignment of stratum $h\in\mathcal{H}=\{1,2,3,4\}$ to $J_{m}^{\lambda}$ ($\square$) or $J_{M}^{\lambda}$ ($\blacksquare$) for a given $\lambda$.
population FPIA optimum allocation
$h$ $A_{h}$ $m_{h}$ $M_{h}$ $\tfrac{A_{h}^{2}}{m_{h}^{2}}$ $\tfrac{A_{h}^{2}}{M_{h}^{2}}$ $J_{m}^{\lambda_{0}}/J_{M}^{\lambda_{0}}$ $J_{m}^{\lambda_{1}}/J_{M}^{\lambda_{1}}$ $J_{m}^{\lambda^{*}}/J_{M}^{\lambda^{*}}$ $\mathbf{x}^{*}$
1 4160 5 50 692224 6922.24 $\blacksquare$ - $44.35$
2 240 5 50 2304 23.04 $\square$ $\square$ $5$
3 530 5 50 11236 112.36 $5.65$
4 40 5 50 64 0.64 $\square$ $\square$ $5$
total sample size $n=60$ $\lambda_{0}=6861.36$ $\lambda_{1}=\tfrac{1}{0}$ $\lambda^{*}=8798.44$
Figure A.1 shows the graphs of the functions $g$ and $\phi$ for the problem considered in this example.
Figure A.1: Functions $g$ and $\phi$ for Example A.2 of the allocation problem for which the FPIA gets blocked.
[…]
We note that a simple remedy to avoid the blocking scenario in the case of a regular allocation is to change $\lambda$ to $\lambda^{new}=\tfrac{1}{\lambda}$, with a corresponding redefinition of all the objects in the original algorithm.
The next Example A.3 illustrates a situation when FPIA does not converge.
Example A.3.
Consider a population given in Table A.3. Initial value of $\lambda_{0}=695.64$ causes lack of convergence of the FPIA due to oscillations:
[…]
Table A.3: Details of FPIA performance for a population with four strata. Here, $\lambda_{0}=\tfrac{1}{s^{2}(\emptyset,\,\emptyset)}$, $\lambda_{k+1}$ is as in (A.6), and $\lambda^{*}=\tfrac{1}{s^{2}(J_{m}^{\lambda^{*}},\,J_{M}^{\lambda^{*}})}$. Squares in cells indicate assignment of stratum $h\in\mathcal{H}=\{1,2,3,4\}$ to $J_{m}^{\lambda}$ ($\square$) or $J_{M}^{\lambda}$ ($\blacksquare$) for a given $\lambda$.
population FPIA optimum allocation
$h$ $A_{h}$ $m_{h}$ $M_{h}$ $\tfrac{A_{h}^{2}}{m_{h}^{2}}$ $\tfrac{A_{h}^{2}}{M_{h}^{2}}$ $J_{m}^{\lambda_{0}}/J_{M}^{\lambda_{0}}$ $J_{m}^{\lambda_{1}}/J_{M}^{\lambda_{1}}$ $J_{m}^{\lambda_{2}}/J_{M}^{\lambda_{2}}$cdots $J_{m}^{\lambda^{*}}/J_{M}^{\lambda^{*}}$ $\mathbf{x}^{*}$
1 380 10 50 1444 57.76 $\square$ $\cdots$ $13.1$
2 140 10 50 196 7.84 $\square$ $\square$ $\square$ $\square$ $10$
3 230 10 50 529 21.16 $\square$ $\square$ $\square$ $\square$ $10$
4 1360 10 50 18496 739.84 $\blacksquare$ $\blacksquare$ $46.9$
total sample size $n=80$ $\lambda_{0}=695.64$ $\lambda_{1}=1444$ $\lambda_{2}=739.84$cdots $\lambda^{*}=841$
Figure A.2 shows the graphs of the functions $g$ and $\phi$ for the problem considered in this example.
Figure A.2: Functions $g$ and $\phi$ for Example A.2 of the allocation problem for which the FPIA does not converge.
The issue of determining a proper starting point $\lambda_{0}$ was considered in MSW, with a recommendation to choose $\lambda_{0}=\tfrac{1}{s^{2}(\emptyset,\,\emptyset)}$. Alternatively, in case when $\tfrac{1}{s^{2}(\emptyset,\,\emptyset)}$ is not close enough to the optimal $\lambda^{*}$ (which is not known a priori), MSW suggests to first run several iterations of a root finding algorithm to get the starting point $\lambda_{0}$ for the FPIA.
Appendix B Appendix: Proof of Theorem 3.1
Remark B.1.
Problem 1.1 is a convex optimization problem as its objective function $f:{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}\to{\mathbb{R}}_{+}$,
$f(\mathbf{x})=\sum_{h\in\mathcal{H}}\tfrac{A_{h}^{2}}{x_{h}},$ (B.1)
and inequality constraint functions $g_{h}^{m}:{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}\to{\mathbb{R}},\,g_{h}^{M}:{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}\to{\mathbb{R}}$,
$g_{h}^{m}(\mathbf{x})=m_{h}-x_{h},\mathrlap{\qquad h\in\mathcal{H},}$ (B.2)
$g_{h}^{M}(\mathbf{x})=x_{h}-M_{h},\mathrlap{\qquad h\in\mathcal{H},}$ (B.3)
are convex functions, whilst the equality constraint function $w:{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}\to{\mathbb{R}}$,
$w(\mathbf{x})=\sum_{h\in\mathcal{H}}x_{h}-n\$
is affine. More specifically, Problem 1.1 is a convex optimization problem of a particular type in which inequality constraint functions (B.2) - (B.3) are affine. See the Appendix E for the definition of the convex optimization problem.
Proof of Theorem 3.1.
[…]
$\nabla f(\mathbf{x})=(-\tfrac{A_{h}^{2}}{x_{h}^{2}},\,h\in\mathcal{H}),\quad\nabla w(\mathbf{x})=\underline{1},\quad\nabla g_{h}^{m}(\mathbf{x})=-\nabla g_{h}^{M}(\mathbf{x})=-\underline{1}_{h},\quad h\in\mathcal{H},$
where, $\underline{1}$ is a vector with all entries $1$ and $\underline{1}_{h}$ is a vector with all entries 00 except the entry with the label $h$, which is $1$. Hence, the KKT conditions (E.2) for Problem 1.1 assume the form
$-\tfrac{A_{h}^{2}}{x^{*2}_{h}}+\lambda-\mu_{h}^{m}+\mu_{h}^{M}=0,\mathrlap{\qquad h\in\mathcal{H},}$ (B.4)
$\sum_{h\in\mathcal{H}}x_{h}^{*}-n=0,$ (B.5)
$m_{h}\leq x^{*}_{h}\leq M_{h},\mathrlap{\qquad h\in\mathcal{H},}$ (B.6)
$\mu^{m}_{h}(m_{h}-x^{*}_{h})=0,\mathrlap{\qquad h\in\mathcal{H},}$ (B.7)
$\mu_{h}^{M}(x^{*}_{h}-M_{h})=0,\mathrlap{\qquad h\in\mathcal{H}.}$ (B.8)
To prove Theorem 3.1, it suffices to show that for $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ with *$\mathcal{L}^{*},\,\mathcal{U}^{*}$ satisfying conditions of CASE I or CASE II, there exist $\lambda\in{\mathbb{R}}$ and $\mu^{m}_{h},\,\mu^{M}_{h}\geq 0,\,h\in\mathcal{H}$, such that (B.4) - (B.8) hold. It should also be noted that the requirement $m_{h}<M_{h},\,h\in\mathcal{H}$, guarantees that $\mathcal{L}^{*}$ and $\mathcal{U}^{*}$ defined in (3.3) and (3.4) are disjoint. Therefore, $\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ is well-defined according to Definition 3.2.
  1. CASE I:
    Take $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ with $\mathcal{L}^{*}$ and $\mathcal{U}^{*}$ as in (3.3). Then, (B.5) is clearly met after referring to (3.2) and (3.1), while (B.6) follows directly from (3.2) and (3.3), since (3.3) for $h\in\mathcal{H}\setminus(\mathcal{L}^{*}\cup\mathcal{U}^{*})$ specifically implies $m_{h}<A_{h}\,s(\mathcal{L}^{*},\,\mathcal{U}^{*})<M_{h}$. Take $\lambda=\tfrac{1}{s^{2}(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ and
    $\mu_{h}^{m}=\begin{cases}\lambda-\tfrac{A_{h}^{2}}{m_{h}^{2}},&h\in\mathcal{L}^{*}\\ 0,&h\in\mathcal{H}\setminus\mathcal{L}^{*},\end{cases}\qquad\qquad\mu_{h}^{M}=\begin{cases}\tfrac{A_{h}^{2}}{M_{h}^{2}}-\lambda,&h\in\mathcal{U}^{*}\\ 0,&h\in\mathcal{H}\setminus\mathcal{U}^{*}.\end{cases}$ (B.9)
    Note that (3.3) along with requirement $n\geq\sum_{h\in\mathcal{H}}m_{h}$ (the latter needed if $\mathcal{U}^{*}=\emptyset$) ensure $s(\mathcal{L}^{*},\,\mathcal{U}^{*})>0$, whilst (3.3) alone implies $\mu^{m}_{h},\,\mu^{M}_{h}\geq 0,\,h\in\mathcal{H}$. After referring to (3.2), it is a matter of simple algebra to verify (B.4), (B.7) and (B.8) for $\lambda,\,\mu^{m}_{h},\,\mu^{M}_{h},\,h\in\mathcal{H}$ defined above.
  2. CASE II:
    Take $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ with *$\mathcal{L}^{*},\,\mathcal{U}^{*}$ satisfying (3.4) and (3.5). Then, condition (B.5) becomes (3.5), while (B.6) is trivially met due to (3.2). Assume that $\mathcal{L}^{*}\neq\emptyset$ and $\mathcal{U}^{*}\neq\emptyset$ (for empty $\mathcal{L}^{*}$ or $\mathcal{U}^{*}$, (B.4), (B.7) and (B.8) are trivially met). Take an arbitrary $\tilde{s}>0$ such that
    $\tilde{s}\in\left[\max_{h\in\mathcal{U}^{*}}\tfrac{M_{h}}{A_{h}},\;\min_{h\in\mathcal{L}^{*}}\tfrac{m_{h}}{A_{h}}\right].$ (B.10)
    Note that (3.4) ensures that the interval above is well-defined. Let $\lambda=\tfrac{1}{\tilde{s}^{2}}$ and $\mu_{h}^{m},\,\mu_{h}^{M},\,h\in\mathcal{H}$ be as in (B.9). Note that (B.10) ensures that $\mu_{h}^{m},\,\mu_{h}^{M}\geq 0$ for all $h\in\mathcal{H}$. Then it is easy to check, similarly as in CASE I, that (B.4), (B.7) and (B.8) are satisfied.
Appendix C Appendix: Auxiliary lemmas and proof of Theorem 4.1
C.1 Notation
Throughout the Appendix C, by $\mathcal{U}_{r},\,\mathcal{L}_{r},\,\widetilde{\mathcal{L}}_{r}$, we denote sets $\mathcal{U},\,\mathcal{L},\,\widetilde{\mathcal{L}}$ respectively, as they are in the $r$-th iteration of RNABOX algorithm after Step 4 and before Step 5. The iteration index $r$ takes on values from set $\{1,\ldots,r^{*}\}$, where $r^{*}\geq 1$ indicates the final iteration of the algorithm. Under this notation, we have $\mathcal{L}_{1}=\emptyset$ and in general, for subsequent iterations, if any (i.e. if $r^{*}\geq 2$), we get
$*.\mathcal{L}_{r}=\mathcal{L}_{r-1}\cup\widetilde{\mathcal{L}}_{r-1}=\bigcup\limits_{i=1}^{r-1}\widetilde{\mathcal{L}}_{i},\mathrlap{\qquad r=2,\ldots,r^{*}.}$ (C.1)
As RNABOX iterates, objects denoted by symbols $n$ and $\mathcal{H}$ are being modified. However, in this Appendix C, whenever we refer to $n$ and $\mathcal{H}$, they always denote the unmodified total sample size and the set of strata labels as in the input of RNABOX. In particular, this is also related to set function $s$ (defined in (3.1)) which depends on $n$ and $\mathcal{H}$.
For convenient notation, for any $\mathcal{A}\subseteq\mathcal{H}$ and any set of real numbers $z_{h},\,h\in\mathcal{A}$, we denote
$z_{\mathcal{A}}=\sum_{h\in\mathcal{A}}z_{h}.$
C.2 Auxiliary remarks and lemmas
We start with a lemma describing important monotonicity properties of function $s$.
Lemma C.1.
Let $\mathcal{A}\subseteq\mathcal{B}\subseteq\mathcal{H}$ and $\mathcal{C}\subseteq\mathcal{D}\subseteq\mathcal{H}$.
  1. 1.
    If $\mathcal{B}\cup\mathcal{D}\subsetneq\mathcal{H}$ and $\mathcal{B}\cap\mathcal{D}=\emptyset$, then
    $s(\mathcal{A},\,\mathcal{C})\geq s(\mathcal{B},\,\mathcal{D})\quad\Leftrightarrow\quad s(\mathcal{A},\,\mathcal{C})(A_{\mathcal{B}\setminus\mathcal{A}}+A_{\mathcal{D}\setminus\mathcal{C}})\leq m_{\mathcal{B}\setminus\mathcal{A}}+M_{\mathcal{D}\setminus\mathcal{C}.}$ (C.2)
  2. 2.
    If $\mathcal{A}\cup\mathcal{D}\subsetneq\mathcal{H}$, $\mathcal{A}\cap\mathcal{D}=\emptyset$, $\mathcal{B}\cup\mathcal{C}\subsetneq\mathcal{H}$, $\mathcal{B}\cap\mathcal{C}=\emptyset$, then
    $s(\mathcal{A},\,\mathcal{D})\geq s(\mathcal{B},\,\mathcal{C})\quad\Leftrightarrow\quad s(\mathcal{A},\,\mathcal{D})(A_{\mathcal{B}\setminus\mathcal{A}}-A_{\mathcal{D}\setminus\mathcal{C}})\leq m_{\mathcal{B}\setminus\mathcal{A}}-M_{\mathcal{D}\setminus\mathcal{C}.}$ (C.3)
Proof.
Clearly, for any $\alpha\in{\mathbb{R}},\,\beta\in{\mathbb{R}},\delta\in{\mathbb{R}}$, $\gamma>0,\,\gamma+\delta>0$, we have
$\tfrac{\alpha+\beta}{\gamma+\delta}\geq\tfrac{\alpha}{\gamma}\quad\Leftrightarrow\quad\tfrac{\alpha+\beta}{\gamma+\delta}\delta\leq\beta.$ (C.4)
To prove (C.2), take
$\alpha$ $=n-m_{\mathcal{B}}-M_{\mathcal{D}}$ $\beta$ $=m_{\mathcal{B}\setminus\mathcal{A}}+M_{\mathcal{D}\setminus\mathcal{C}}$
$\gamma$ $=A_{\mathcal{H}}-A_{\mathcal{B}\cup\mathcal{D}}$ $\delta$ $=A_{\mathcal{B}\setminus\mathcal{A}}+A_{\mathcal{D}\setminus\mathcal{C}}.$
Then, $\tfrac{\alpha}{\gamma}=s(\mathcal{B},\,\mathcal{D}),\,\tfrac{\alpha+\beta}{\gamma+\delta}=s(\mathcal{A},\,\mathcal{C})$, and hence (C.2) holds as an immediate consequence of (C.4).
Similarly for (C.3), take
$\alpha$ $=n-m_{\mathcal{B}}-M_{\mathcal{C}}$ $\beta$ $=m_{\mathcal{B}\setminus\mathcal{A}}-M_{\mathcal{D}\setminus\mathcal{C}}$
$\gamma$ $=A_{\mathcal{H}}-A_{\mathcal{B}\cup\mathcal{C}}$ $\delta$ $=A_{\mathcal{B}\setminus\mathcal{A}}-A_{\mathcal{D}\setminus\mathcal{C}},$
and note that $\gamma+\delta=A_{\mathcal{H}}-A_{\mathcal{B}\cup\mathcal{C}}+A_{\mathcal{B}\setminus\mathcal{A}}-A_{\mathcal{D}\setminus\mathcal{C}}=A_{\mathcal{H}}-A_{\mathcal{B}}-A_{\mathcal{C}}+A_{\mathcal{B}}-A_{\mathcal{A}}-A_{\mathcal{D}}+A_{\mathcal{C}}=A_{\mathcal{H}}-A_{\mathcal{A}\cup\mathcal{D}}>0$ due to the assumptions made for $\mathcal{A},\,\mathcal{D},\,\mathcal{B},\,\mathcal{C}$, and $A_{h}>0,\,h\in\mathcal{H}$. Then, $\tfrac{\alpha}{\gamma}=s(\mathcal{B},\,\mathcal{C})$, $\tfrac{\alpha+\beta}{\gamma+\delta}=s(\mathcal{A},\,\mathcal{D})$, and hence (C.3) holds as an immediate consequence of (C.4). ∎
The remark below describes some relations between sets $\mathcal{L}_{r}$ and $\mathcal{U}_{r},\,r=1,\ldots,r^{*}\geq 1$, appearing in RNABOX algorithm. These relations are particularly important for understanding computations involving the set function $s$ (recall, that it is defined only for such two disjoint sets, the union of which is a proper subset of $\mathcal{H}$).
Remark C.1.
For $r^{*}\geq 1$,
$\mathcal{L}_{r}\cap\mathcal{U}_{r}=\emptyset,\mathrlap{\qquad r=1,\ldots,r^{*},}$ (C.5)
and for $r^{*}\geq 2$,
$\mathcal{L}_{r}\cup\mathcal{U}_{r}\subsetneq\mathcal{H},\mathrlap{\qquad r=1,\ldots,r^{*}-1.}$ (C.6)
Moreover, let $\mathbf{x}^{*}$ be as in Step 6 of RNABOX algorithm. Then, for $r^{*}\geq 1$,
[…]
and
[…]
Proof.
From the definition of set $\mathcal{U}$ in Step 3 of RNABOX, for $r^{*}\geq 1$,
$\mathcal{U}_{r}\subseteq\mathcal{H}\setminus\mathcal{L}_{r},\mathrlap{\qquad r=1,\ldots,r^{*},}$ (C.9)
which proves (C.5). Following (C.1), for $r^{*}\geq 2$,
$\mathcal{L}_{r}=\bigcup\limits_{i=1}^{r-1}\widetilde{\mathcal{L}}_{i}\subseteq\mathcal{H},\mathrlap{\qquad r=2,\ldots,r^{*},}$ (C.10)
where the inclusion is due to definition of set $\widetilde{\mathcal{L}}$ in Step 4 of RNABOX, i.e. ~$\widetilde{\mathcal{L}}_{r}\subseteq\mathcal{H}\setminus(\mathcal{L}_{r}\cup\mathcal{U}_{r})$ for $r=1,\ldots,r^{*}$. Inclusions (C.9), (C.10) with $\mathcal{L}_{1}=\emptyset$ imply
$\mathcal{L}_{r}\cup\mathcal{U}_{r}\subseteq\mathcal{H},\mathrlap{\qquad r=1,\ldots,r^{*}\geq 1.}$ (C.11)
Given that $r^{*}\geq 2$, Step 5 of the algorithm ensures that set ~$\widetilde{\mathcal{L}}_{r}\subseteq\mathcal{H}\setminus(\mathcal{L}_{r}\cup\mathcal{U}_{r})$ is non-empty for $r=1,\ldots,r^{*}-1$, which implies $\mathcal{L}_{r}\cup\mathcal{U}_{r}\neq\mathcal{H}$. This fact combined with (C.11) gives (C.6). Equivalences (C.7) and (C.8) hold trivially after referring to Definition 1.1 of regular and vertex allocations. ∎
The following two remarks summarize some important facts arising from Step 3 of RNABOX algorithm. These facts will serve as starting points for most of the proofs presented in this section.
Remark C.2.
In each iteration $r=1,\ldots,r^{*}\geq 1$, of RNABOX algorithm, a vector $(x^{**}_{h},\,h\in\mathcal{H}\setminus\mathcal{L}_{r})$ obtained in Step 3, has the elements of the form
[…]
where the set function $s$ is defined in (3.1). Equation (C.12) is a direct consequence of Theorem D.1.
Remark C.3.
Remark C.2 together with Theorem D.1, for $r^{*}\geq 2$ yield
$\mathcal{U}_{r}=\{h\in\mathcal{H}\setminus\mathcal{L}_{r}:\,A_{h}\,s(\mathcal{L}_{r},\,\mathcal{U}_{r})\geq M_{h}\},\mathrlap{\qquad r=1,\ldots,r^{*}-1,}$ (C.13)
whilst for $r^{*}\geq 1$,
$\mathcal{U}_{r^{*}}=\{h\in\mathcal{H}\setminus\mathcal{L}_{r^{*}}:\,A_{h}\,s(\mathcal{L}_{r^{*}},\,\mathcal{U}_{r^{*}})\geq M_{h}\},$ (C.14)
if and only if $\mathbf{x}^{*}$ (computed at Step 6 of RNABOX algorithm) is: a regular allocation or a vertex allocation with $\mathcal{L}_{r^{*}}=\mathcal{H}$. Moreover, for $r^{*}\geq 1$,
~$*.\widetilde{\mathcal{L}}_{r}=\{h\in\mathcal{H}\setminus(\mathcal{L}_{r}\cup\mathcal{U}_{r}):\,A_{h}\,s(\mathcal{L}_{r},\,\mathcal{U}_{r})\leq m_{h}\},\mathrlap{\qquad r=1,\ldots,r^{*}.}$ (C.15)
Note that in Remark C.3, function $s$ is well-defined due to Remark C.1. The need to limit the scope of (C.14) to regular allocations only, is dictated by the fact that in the case of a vertex allocation we have $\mathcal{L}_{r^{*}}\cup\mathcal{U}_{r^{*}}=\mathcal{H}$ (see (C.8)) and therefore $s(\mathcal{L}_{r^{*}},\,\mathcal{U}_{r^{*}})$ is not well-defined.
Lemma C.2 and Lemma C.3 reveal certain monotonicity properties of sequence $(\mathcal{U}_{r})_{r=1}^{r^{*}}$ and sequence $\left(s(\mathcal{L}_{r},\,\mathcal{U}_{r})\right)_{r=1}^{r^{*}}$, respectively. These properties will play a crucial role in proving Theorem 4.1.
Lemma C.2.
Sequence $(\mathcal{U}_{r})_{r=1}^{r^{*}}$ is non-increasing, that is, for $r^{*}\geq 2$,
$\mathcal{U}_{r}\supseteq\mathcal{U}_{r+1},\mathrlap{\qquad r=1,\ldots,r^{*}-1.}$ (C.16)
Proof.
Let $r^{*}\geq 2$ and $r=1,\ldots,r^{*}-1$. Then, by (C.6), $\mathcal{L}_{r}\cup\mathcal{U}_{r}\subsetneq\mathcal{H}$. Following (C.13), the domain of discourse for $\mathcal{U}_{r}$ is $\mathcal{H}\setminus\mathcal{L}_{r}$, and in fact it is $\mathcal{H}\setminus(\mathcal{L}_{r}\cup\widetilde{\mathcal{L}}_{r})=\mathcal{H}\setminus\mathcal{L}_{r+1}$, since $\mathcal{U}_{r}\not\subset\widetilde{\mathcal{L}}_{r}$ as ensured by Step 4 of RNABOX. That is, both $\mathcal{U}_{r}$ and $\mathcal{U}_{r+1}$ have essentially the same domain of discourse, which is $\mathcal{H}\setminus\mathcal{L}_{r+1}$. Given this fact and the form of the set-builder predicate in (C.13) - (C.14) as well as equality $\mathcal{U}_{r^{*}}=\mathcal{H}\setminus\mathcal{L}_{r^{*}}$ for the case when $\mathbf{x}^{*}$ is a vertex allocation (for which (C.14) does not apply), we conclude that only one of the following two distinct cases is possible: $\mathcal{U}_{r}\supseteq\mathcal{U}_{r+1}$ or $\mathcal{U}_{r}\subsetneq\mathcal{U}_{r+1}$. The proof is by contradiction, that is, assume that (C.16) does not hold. Therefore, in view of the above observation, there exists $r\in\{1,\ldots,r^{*}-1\}$ such that $\mathcal{U}_{r}\subsetneq\mathcal{U}_{r+1}$. Then,
$\emptyset\neq(\mathcal{U}_{r+1}\setminus\mathcal{U}_{r})\subsetneq\mathcal{H}\setminus(\mathcal{L}_{r}\cup\mathcal{U}_{r}),$ (C.17)
and hence, due to (C.12),
$A_{h}\,s(\mathcal{L}_{r},\,\mathcal{U}_{r})<M_{h},\mathrlap{\qquad h\in\mathcal{U}_{r+1}\setminus\mathcal{U}_{r}.}$ (C.18)
On the other hand, from (C.15),
$A_{h}\,s(\mathcal{L}_{r},\,\mathcal{U}_{r})\leq m_{h},\mathrlap{\qquad h\in\widetilde{\mathcal{L}}_{r}.}$ (C.19)
Summing sidewise: (C.18) over $h\in\mathcal{U}_{r+1}\setminus\mathcal{U}_{r}$, (C.19) over $h\in\widetilde{\mathcal{L}}_{r}$, and then all together, we get
$s(\mathcal{L}_{r},\,\mathcal{U}_{r})(A_{\widetilde{\mathcal{L}}_{r}}+A_{\mathcal{U}_{r+1}\setminus\mathcal{U}_{r}})<m_{\widetilde{\mathcal{L}}_{r}}+M_{\mathcal{U}_{r+1}\setminus\mathcal{U}_{r}}.$ (C.20)
  1. Vector $\mathbf{x}^{*}$ is a regular allocation:
    In this case, following Remark C.1, we see that inequality (C.20) is the right-hand side of equivalence (C.2) with
    $\mathcal{A}$ $=\mathcal{L}_{r}\subseteq(\mathcal{L}_{r}\cup\widetilde{\mathcal{L}}_{r})=\mathcal{L}_{r+1}=\mathcal{B}\subsetneq\mathcal{H},$ (C.21)
    $\mathcal{C}$ $=\mathcal{U}_{r}\subsetneq\mathcal{U}_{r+1}=\mathcal{D}\subsetneq\mathcal{H}.$ (C.22)
    Then, following Lemma C.1, inequality (C.20) is equivalent to
    $s(\mathcal{L}_{r},\mathcal{U}_{r})>s(\mathcal{L}_{r+1},\mathcal{U}_{r+1}).$ (C.23)
    Combining
    $s(\mathcal{L}_{r+1},\,\mathcal{U}_{r+1})\geq\tfrac{M_{h}}{A_{h}},\mathrlap{\qquad h\in\mathcal{U}_{r+1},}$ (C.24)
    (it follows from (C.13)-(C.14)) with inequalities (C.23) and (C.18), we get the contradiction
    $\tfrac{M_{h}}{A_{h}}>s(\mathcal{L}_{r},\,\mathcal{U}_{r})>s(\mathcal{L}_{r+1},\,\mathcal{U}_{r+1})\geq\tfrac{M_{h}}{A_{h}},\mathrlap{\qquad h\in\mathcal{U}_{r+1}\setminus\mathcal{U}_{r}.}$ (C.25)
    Therefore, (C.16) holds true, given that $\mathcal{L}_{r^{*}}\cup\mathcal{U}_{r^{*}}\subsetneq\mathcal{H}$.
  2. Vector $\mathbf{x}^{*}$ is a vertex allocation:
    Since $\mathcal{L}_{r+1}\cup\mathcal{U}_{r+1}\subsetneq\mathcal{H}$ for $r=1,\ldots,r^{*}-2$, the proof of (C.16) for such $r$ is identical to the proof for the case of regular allocation. Hence, we only need to show that (C.16) holds for $r=r^{*}-1$. For this purpose, we will exploit inequality (C.20), which in view of Definition 3.1 of set function $s$, assumes the following form for $r=r^{*}-1$,
    $\tfrac{n-m_{\mathcal{L}_{r^{*}-1}}-M_{\mathcal{U}_{r^{*}-1}}}{A_{\mathcal{H}}-A_{\mathcal{L}_{r^{*}-1}\cup\mathcal{U}_{r^{*}-1}}}\,(A_{\widetilde{\mathcal{L}}_{r^{*}-1}}+A_{\mathcal{U}_{r^{*}}\setminus\mathcal{U}_{r^{*}-1}})<m_{\widetilde{\mathcal{L}}_{r^{*}-1}}+M_{\mathcal{U}_{r^{*}}\setminus\mathcal{U}_{r^{*}-1}}.$ (C.26)
    Since $A_{\mathcal{H}}-A_{\mathcal{L}_{r^{*}-1}\cup\mathcal{U}_{r^{*}-1}}=A_{\widetilde{\mathcal{L}}_{r^{*}-1}}+A_{\mathcal{U}_{r^{*}}\setminus\mathcal{U}_{r^{*}-1}}$, for $\mathcal{L}_{r^{*}}\cup\mathcal{U}_{r^{*}}=\mathcal{H}$, inequality (C.26) simplifies to
    $n<m_{\widetilde{\mathcal{L}}_{r^{*}-1}}+m_{\mathcal{L}_{r^{*}-1}}+M_{\mathcal{U}_{r^{*}}\setminus\mathcal{U}_{r^{*}-1}}+M_{\mathcal{U}_{r^{*}-1}}=m_{\mathcal{L}_{r^{*}}}+M_{\mathcal{U}_{r^{*}}}=n,$ (C.27)
    which is a contradiction. Note that the last equality follows from Step 3 of the RNABOX after referring to (D.3) and using the fact that $\mathcal{U}_{r^{*}}=\mathcal{H}\setminus\mathcal{L}_{r^{*}}$ for a vertex allocation. Therefore, (C.16) holds true also for $\mathcal{L}_{r^{*}}\cup\mathcal{U}_{r^{*}}=\mathcal{H}$.
Lemma C.3.
Let $r^{*}\geq 3$. Then
$s(\mathcal{L}_{r},\,\mathcal{U}_{r})\geq s(\mathcal{L}_{r+1},\,\mathcal{U}_{r+1}),\mathrlap{\qquad r=1,\ldots,r^{*}-2.}$ (C.28)
Moreover, if $\mathbf{x}^{*}$ (computed at Step 6 of RNABOX algorithm) is a regular allocation and $r^{*}\geq 2$, then
$s(\mathcal{L}_{r^{*}-1},\,\mathcal{U}_{r^{*}-1})\geq s(\mathcal{L}_{r^{*}},\,\mathcal{U}_{r^{*}}).$ (C.29)
Proof.
We first prove (C.28). Let $r^{*}\geq 3$ and $r=1,\ldots,r^{*}-2$. Following Lemma C.2 and using (C.13),
$A_{h}\,s(\mathcal{L}_{r},\,\mathcal{U}_{r})\geq M_{h},\mathrlap{\qquad h\in\mathcal{U}_{r}\setminus\mathcal{U}_{r+1}.}$ (C.30)
On the other hand, from (C.15),
$A_{h}\,s(\mathcal{L}_{r},\,\mathcal{U}_{r})\leq m_{h},\mathrlap{\qquad h\in\widetilde{\mathcal{L}}_{r}.}$ (C.31)
Multiplying both sides of inequality (C.30) by $-1$, summing it sidewise over $h\in\mathcal{U}_{r}\setminus\mathcal{U}_{r+1}$ and then adding it to (C.31), which is previously summed sidewise over $h\in\widetilde{\mathcal{L}}_{r}$, we get
$s(\mathcal{L}_{r},\,\mathcal{U}_{r})(A_{\widetilde{\mathcal{L}}_{r}}-A_{\mathcal{U}_{r}\setminus\mathcal{U}_{r+1}})\leq m_{\widetilde{\mathcal{L}}_{r}}-M_{\mathcal{U}_{r}\setminus\mathcal{U}_{r+1}}.$ (C.32)
Relation (C.32) is the second inequality in (C.3) with
$\mathcal{A}$ $=\mathcal{L}_{r}\subsetneq(\mathcal{L}_{r}\cup\widetilde{\mathcal{L}}_{r})=\mathcal{L}_{r+1}=\mathcal{B}\subsetneq\mathcal{H},$ (C.33)
$\mathcal{C}$ $=\mathcal{U}_{r+1}\subseteq\mathcal{U}_{r}=\mathcal{D}\subsetneq\mathcal{H}.$ (C.34)
Based on Remark C.1, we see that $\mathcal{A}\cup\mathcal{D}\subsetneq\mathcal{H}$, $\mathcal{A}\cap\mathcal{D}=\emptyset$, and $\mathcal{B}\cup\mathcal{C}\subsetneq\mathcal{H}$, $\mathcal{B}\cap\mathcal{C}=\emptyset$, and thus the first inequality in (C.3) follows, that is
$s(\mathcal{L}_{r},\,\mathcal{U}_{r})\geq s(\mathcal{L}_{r+1},\,\mathcal{U}_{r+1}).$ (C.35)
Hence (C.28) is proved. If $\mathbf{x}^{*}$ is a regular allocation, in view of Remark C.1, the same reasoning leading to inequality (C.35) clearly remains valid for $r=r^{*}-1,\,r^{*}\geq 2$. ∎
C.3 Proof of Theorem 4.1
To prove Theorem 4.1, we have to show that:
  1. (I)
    the algorithm terminates in a finite number of iterations, i.e. $r^{*}<\infty$,
  2. (II)
    the solution computed at $r^{*}$ is optimal.
The proof of part (I) is relatively straightforward. In every iteration $r=1,\ldots,r^{*}-1,\,r^{*}\geq 2$, the set of strata labels $\mathcal{H}$ is reduced by subtracting ~$\widetilde{\mathcal{L}}_{r}$. Therefore, $r^{*}\leq\lvert\mathcal{H}\rvert+1<\infty$, where $r^{*}=\lvert\mathcal{H}\rvert+1$ if and only if |~$\lvert\widetilde{\mathcal{L}}_{r}\rvert=1,\,r=1,\ldots,r^{*}-1$. In words, the algorithm stops after at most $\lvert\mathcal{H}\rvert+1$ iterations.
In order to prove part (II), following Theorem 3.1 and Remark 3.1, it suffices to show that when $\mathbf{x}^{*}$ (computed at Step 6 of RNABOX algorithm) is a regular allocation, for all $h\in\mathcal{H}$,
$h\in\mathcal{L}_{r^{*}}\quad\Leftrightarrow\quad s(\mathcal{L}_{r^{*}},\,\mathcal{U}_{r^{*}})\leq\tfrac{m_{h}}{A_{h}},$ (C.36)
$h\in\mathcal{U}_{r^{*}}\quad\Leftrightarrow\quad s(\mathcal{L}_{r^{*}},\,\mathcal{U}_{r^{*}})\geq\tfrac{M_{h}}{A_{h}},$ (C.37)
and when $\mathbf{x}^{*}$ is a vertex allocation
$\max_{h\in\mathcal{U}_{r^{*}}}\tfrac{M_{h}}{A_{h}}\leq\min_{h\in\mathcal{L}_{r^{*}}}\tfrac{m_{h}}{A_{h}},\quad\text{when }\mathcal{U}_{r^{*}}\neq\emptyset\text{ and }\mathcal{L}_{r^{*}}\neq\emptyset,$ (C.38)
$m_{\mathcal{L}_{r^{*}}}+M_{\mathcal{U}_{r^{*}}}=n.$ (C.39)
  1. Vector $\mathbf{x}^{*}$ is a regular allocation:
    Note that Remark C.1, implies that $s(\mathcal{L}_{r},\,\mathcal{U}_{r})$ is well-defined. We start with equivalence (C.36).
    Necessity: For $r^{*}=1$, we have $\mathcal{L}_{r^{*}}=\emptyset$ and hence, the right-hand side of equivalence (C.36) is trivially met. Let $r^{*}\geq 2$, and $h\in\mathcal{L}_{r^{*}}=\bigcup\limits_{r=1}^{r^{*}-1}\widetilde{\mathcal{L}}_{r}$. Thus, $h\in\widetilde{\mathcal{L}}_{r}$ for some $r\in\{1,\ldots,r^{*}-1\}$ and then, due to (C.15), we have $s(\mathcal{L}_{r},\,\mathcal{U}_{r})\leq\tfrac{m_{h}}{A_{h}}$. Consequently, (C.28) with (C.29) yield $s(\mathcal{L}_{r^{*}},\,\mathcal{U}_{r^{*}})\leq\tfrac{m_{h}}{A_{h}}$.
    Sufficiency: Since ~$\widetilde{\mathcal{L}}_{r^{*}}=\emptyset$, (C.15) implies
    $\{h\in\mathcal{H}\setminus(\mathcal{L}_{r^{*}}\cup\mathcal{U}_{r^{*}}):\,s(\mathcal{L}_{r^{*}},\,\mathcal{U}_{r^{*}})\leq\tfrac{m_{h}}{A_{h}}\}=\emptyset,\mathrlap{\qquad r^{*}\geq 1.}$ (C.40)
    On the other hand, (C.14) along with $m_{h}<M_{h},\,h\in\mathcal{H}$, yield
    $s(\mathcal{L}_{r^{*}},\,\mathcal{U}_{r^{*}})\geq\tfrac{M_{h}}{A_{h}}>\tfrac{m_{h}}{A_{h}},\mathrlap{\qquad h\in\mathcal{U}_{r^{*}},}$ (C.41)
    and hence, (C.40) reads
    $\{h\in\mathcal{H}\setminus\mathcal{L}_{r^{*}}:\,s(\mathcal{L}_{r^{*}},\,\mathcal{U}_{r^{*}})\leq\tfrac{m_{h}}{A_{h}}\}=\emptyset,\mathrlap{\qquad r^{*}\geq 1.}$ (C.42)
    The proof of necessity in (C.37) is immediate in view of (C.14), whilst sufficiency follows by contradiction. Indeed, let $r^{*}\geq 1$. Assume that the right-hand side of equivalence (C.37) holds and $h\not\in\mathcal{U}_{r^{*}}$. Then, in view of Remark C.1, either $h\in\mathcal{H}\setminus(\mathcal{L}_{r^{*}}\cup\mathcal{U}_{r^{*}})$ and then from (C.12)
    $s(\mathcal{L}_{r^{*}},\,\mathcal{U}_{r^{*}})<\tfrac{M_{h}}{A_{h}},$ (C.43)
    a contradiction, or $h\in\mathcal{L}_{r^{*}}$ and then from (C.36), in view of $m_{h}<M_{h},\,h\in\mathcal{H}$,
    $s(\mathcal{L}_{r^{*}},\,\mathcal{U}_{r^{*}})\leq\tfrac{m_{h}}{A_{h}}<\tfrac{M_{h}}{A_{h}},$ (C.44)
    a contradiction.
  2. Vector $\mathbf{x}^{*}$ is a vertex allocation:
    For $r^{*}=1$, the only possibility is that $\mathcal{U}_{r^{*}}=\mathcal{H},\,\mathcal{L}_{r^{*}}=\emptyset$. Then, (C.38) is clearly met, while (C.39) follows from Step 3 of the RNABOX after referring to (D.3). Let $r^{*}\geq 2$. Then, by (C.13) we have
    $s(\mathcal{L}_{r^{*}-1},\,\mathcal{U}_{r^{*}-1})\geq\tfrac{M_{h}}{A_{h}},\mathrlap{\qquad h\in\mathcal{U}_{r^{*}-1}\supseteq\mathcal{U}_{r^{*}},}$ (C.45)
    where the set inclusion is due to Lemma C.2. On the other hand, from (C.15), we get
    $s(\mathcal{L}_{r^{*}-1},\,\mathcal{U}_{r^{*}-1})\leq\tfrac{m_{h}}{A_{h}},\mathrlap{\qquad h\in\mathcal{L}_{r^{*}-1}\cup\widetilde{\mathcal{L}}_{r^{*}-1}=\mathcal{L}_{r^{*}},}$ (C.46)
    where the fact that above inequality is met for $h\in\mathcal{L}_{r^{*}-1}$ follows from (C.28). By comparing (C.45) and (C.46) we clearly see that (C.38) is satisfied. Lastly, equation (C.39) is fulfilled due to
    $n-m_{\widetilde{\mathcal{L}}_{1}}-\ldots-m_{\widetilde{\mathcal{L}}_{r*-1}}=n-m_{\mathcal{L}_{r^{*}}}=M_{\mathcal{U}_{r^{*}}},$ (C.47)
    where first equality follows from (C.1) while the second one follows from Step 3 of the RNABOX after referring to (D.3) and using the fact that $\mathcal{U}_{r^{*}}=\mathcal{H}\setminus\mathcal{L}_{r^{*}}$ for a vertex allocation.
Appendix D Appendix: Optimality conditions for Problem 2.1
The following Theorem D.1 provides necessary and sufficient conditions for the optimal solution to Problem 2.1. It was originally given as Theorem 1.1 in Wesołowski et al. (2022) and it is crucial for the proof of Theorem 4.1. Here, we will quote it in a slightly expanded form so that it also covers the case of $\mathcal{U}^{*}=\mathcal{H}$. As usual, the set function $s$ is defined as in Definition 3.1. The algorithm that solves Problem 2.1 is RNA and it is given in Section 4 of this paper.
Theorem D.1.
The optimization Problem 2.1 has a unique optimal solution. Point $\mathbf{x}^{*}=(x_{h}^{*},\,h\in\mathcal{H})\in{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}$ is a solution to optimization Problem 2.1 if and only if $\mathbf{x}^{*}$ has entries of the form
[…]
with $\mathcal{U}^{*}\subseteq\mathcal{H}$, such that one of the following two cases holds:
  1. CASE I:
    $\mathcal{U}^{*}\subsetneq\mathcal{H}$ and
    $\mathcal{U}^{*}=\left\{h\in\mathcal{H}:\,A_{h}\,s(\emptyset,\,\mathcal{U}^{*})\geq M_{h}\right\}.$ (D.2)
  2. CASE II:
    $\mathcal{U}^{*}=\mathcal{H}$ and
    $n=\sum_{h\in\mathcal{H}}M_{h}.$ (D.3)
Appendix E Appendix: Convex optimization scheme and the KKT conditions
A convex optimization problem is an optimization problem in which the objective function is a convex function and the feasible set is a convex set. In standard form it is written as
$\begin{split}\underset{\mathbf{x}\,\in\,\mathcal{D}}{\mathrm{minimize\leavevmode\nobreak\ \,}}&\quad f(\mathbf{x})\\ \mathrm{subject\leavevmode\nobreak\ to}&\quad w_{i}(\mathbf{x})=0,\quad{i=1,\ldots,k}\\ &\quad g_{j}(\mathbf{x})\leq 0,\quad{j=1,\ldots,\ell,}\end{split}$ (E.1)
where $\mathcal{D}\subseteq{\mathbb{R}}^{p},\,p\in{\mathbb{N}}_{+}$, is the optimization variable, the objective function $f:\mathcal{D}_{f}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}}$ and inequality constraint functions $g_{j}:\mathcal{D}_{g_{j}}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}},\,j=1,\ldots,\ell$, are convex, whilst equality constraint functions $w_{i}:\mathcal{D}_{w_{i}}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}},\,i=1,\ldots,k$, are affine. Here, $\mathcal{D}=\mathcal{D}_{f}\cap\bigcap_{i=1}^{k}\mathcal{D}_{w_{i}}\cap\bigcap_{j=1}^{\ell}\mathcal{D}_{g_{j}}$ denotes a common domain of all the functions. Point $\mathbf{x}\in\mathcal{D}$ is called feasible if it satisfies all of the constraints, otherwise the point is called infeasible. An optimization problem is called feasible if there exists $\mathbf{x}\in\mathcal{D}$ that is feasible, otherwise the problem is called infeasible.
In the context of the optimum allocation Problem 1.1 discussed in this paper, we are interested in a particular type of the convex problem, i.e. (E.1) in which all inequality constraint functions $g_{j},\,j=1,\ldots,\ell$, are affine. It is well known, see, e.g. the monograph Boyd and Vandenberghe (2004), that the solution for such an optimization problem can be identified through the set of equations and inequalities known as the Karush-Kuhn-Tucker (KKT) conditions, which in this case are not only necessary but also sufficient.
Theorem E.1 (KKT conditions for convex optimization problem with affine inequality constraints).
A point $\mathbf{x}^{*}\in\mathcal{D}\subseteq{\mathbb{R}}^{p}$ is a solution to the convex optimization problem (E.1) in which functions $g_{j},\,j=1,\ldots,\ell$, are affine if and only if there exist numbers $\lambda_{i}\in{\mathbb{R}}$, $i=1,\ldots,k$, and $\mu_{j}\geq 0$, $j=1,\ldots,\ell$, called KKT multipliers, such that
$\begin{gathered}\nabla f(\mathbf{x}^{*})+\sum_{i=1}^{k}\lambda_{i}\nabla w_{i}(\mathbf{x}^{*})+\sum_{j=1}^{\ell}\mu_{j}\nabla g_{j}(\mathbf{x}^{*})=\mathbf{0}\\ w_{i}(\mathbf{x}^{*})=0,\mathrlap{\qquad i=1,\ldots,k}\\ g_{j}(\mathbf{x}^{*})\leq 0\mathrlap{\qquad j=1,\ldots,\ell}\\ \mu_{j}g_{j}(\mathbf{x}^{*})=0\mathrlap{\qquad j=1,\ldots,\ell.}\end{gathered}$ (E.2)

### v4

@@
 We implement RNABOX in R as a part of our package stratallo which is available from the Comprehensive R Archive Network (CRAN) repository.
-Key Words: Stratified sampling; Optimum sample allocation; Optimum allocation; Optimum allocation under box constraints; Neyman allocation; Recursive Neyman algorithm.
+Key Words: Stratified sampling; Optimum allocation; Optimum allocation under box constraints; Neyman allocation; Recursive Neyman algorithm.
 1 Introduction
@@
 An allocation $\mathbf{x}=(x_{h},\,h\in\mathcal{H})$ is called a vertex one if and only if
-[…]
+$x_{h}=\begin{cases}m_{h},&\mathrlap{\qquad h\in\mathcal{L}}\\ M_{h},&\mathrlap{\qquad h\in\mathcal{U},}\end{cases}$
 where $\mathcal{L},\,\mathcal{U}\subseteq\mathcal{H}$ are such that $\mathcal{L}\cup\mathcal{U}=\mathcal{H}$ and $\mathcal{L}\cap\mathcal{U}=\emptyset$.
@@
 In spite of this fact, a thorough analysis of the literature shows that Problem 1.1 has not been completely understood yet and it suffers from the lack of fully satisfactory algorithms.
-Below, we briefly review existing methods for solving Problem 1.1, including also methods that provide integer-valued solutions.
+Below, we briefly review existing methods for solving Problem 1.1, including methods that provide integer-valued solutions.
 2.1 Not-necessarily integer-valued allocation
@@
 These methods have been involved in the problem of optimum sample allocation since solving the allocation problem is equivalent to finding the extreme (namely, stationary points) of a certain objective function over a feasible set.
-Knowing the (approximate) extreme of the objective function, one can give the (approximate, yet typically sufficiently accurate) sizes of the samples allocated to individual strata.
+Knowing the (approximate) extreme of the objective function, one can determine the (approximate, yet typically sufficiently accurate) sizes of samples allocated to individual strata.
 In a similar yet different approach adopted e.g.
@@
 (2015).
-The proposed procedure that solves that problem, uses binary integer programming algorithm and can be applied to the univariate case.
+The proposed procedure that solves that problem uses binary integer programming algorithm and can be applied to the univariate case.
 See also Brito et al.
@@
 application to the German census in Burgard and Münnich (2012), and it becomes even more pronounced in iterative solutions to stratification problems, when the number of iterations may count in millions (see, e.g.
-Lednicki and Wieczorkowski, 2003, Gunning and Horgan, 2004, Ahmad et al., 2008, Baillargeon and Rivest, 2011, Barcaroli, 2014).
+Lednicki and Wieczorkowski, 2003, Gunning and Horgan, 2004, Khan et al., 2008, Baillargeon and Rivest, 2011, Barcaroli, 2014).
 Having all that said, the search for a new, universal, theoretically sound and computationally effective recursive algorithms of optimum sample allocation under two-sided constraints on the sample strata-sizes, is crucial both for theory and practice of survey sampling.
@@
 The set function $s$ is defined as
-$s(\mathcal{L},\,\mathcal{U})=\frac{n-\sum_{h\in\mathcal{L}}m_{h}-\sum_{h\in\mathcal{U}}M_{h}}{\sum_{h\in\mathcal{H}}A_{h}-\sum_{h\in\mathcal{L}\cup\mathcal{U}}A_{h}}.$ (3.1)
+$s(\mathcal{L},\,\mathcal{U})=\frac{n-\sum_{h\in\mathcal{L}}m_{h}-\sum_{h\in\mathcal{U}}M_{h}}{\sum_{h\in\mathcal{H}\setminus(\mathcal{L}\cup\mathcal{U})}A_{h}}.$ (3.1)
 Below, we will introduce the $\mathbf{x}^{(\mathcal{L},\,\mathcal{U})}$ vector for disjoint $\mathcal{L},\,\mathcal{U}\subseteq\mathcal{H}$.
@@
 We define the vector $\mathbf{x}^{(\mathcal{L},\,\mathcal{U})}=(x_{h}^{(\mathcal{L},\,\mathcal{U})},\,h\in\mathcal{H})$ as follows
-[…]
+$x_{h}^{(\mathcal{L},\,\mathcal{U})}=\begin{cases}m_{h},&\mathrlap{\qquad h\in\mathcal{L}}\\ M_{h},&\mathrlap{\qquad h\in\mathcal{U}}\\ A_{h}\,s(\mathcal{L},\,\mathcal{U}),&\mathrlap{\qquad h\in\mathcal{H}\setminus(\mathcal{L}\cup\mathcal{U}).}\end{cases}$ (3.2)
 The following Theorem 3.1 characterizes the form of the optimal solution to Problem 1.1 and therefore is one of the key results of this paper.
@@
 The optimization Problem 1.1 has a unique optimal solution.
-Point $\mathbf{x}^{*}\in{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}$ is a solution to optimization Problem 1.1 if and only if $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$, with disjoint *$\mathcal{L}^{*},\,\mathcal{U}^{*}\subseteq\mathcal{H}$, such that one of the following two cases holds:
+Point $\mathbf{x}^{*}\in{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}$ is a solution to optimization Problem 1.1 if and only if $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$, with disjoint $\mathcal{L}^{*},\,\mathcal{U}^{*}\subseteq\mathcal{H}$, such that one of the following two cases holds:
   1. CASE I:
-    *$\mathcal{L}^{*}\cup\mathcal{U}^{*}\subsetneq\mathcal{H}$ and
-    *=$\begin{gathered}\mathcal{L}^{*}=\left\{h\in\mathcal{H}:\,s(\mathcal{L}^{*},\,\mathcal{U}^{*})\leq\tfrac{m_{h}}{A_{h}}\right\},\\ \mathcal{U}^{*}=\left\{h\in\mathcal{H}:\,s(\mathcal{L}^{*},\,\mathcal{U}^{*})\geq\tfrac{M_{h}}{A_{h}}\right\}.\end{gathered}$ (3.3)
+    $\mathcal{L}^{*}\cup\mathcal{U}^{*}\subsetneq\mathcal{H}$ and
+    =$\begin{gathered}\mathcal{L}^{*}=\left\{h\in\mathcal{H}:\,s(\mathcal{L}^{*},\,\mathcal{U}^{*})\leq\tfrac{m_{h}}{A_{h}}\right\},\\ \mathcal{U}^{*}=\left\{h\in\mathcal{H}:\,s(\mathcal{L}^{*},\,\mathcal{U}^{*})\geq\tfrac{M_{h}}{A_{h}}\right\}.\end{gathered}$ (3.3)
   2. CASE II:
-    *$\mathcal{L}^{*}\cup\mathcal{U}^{*}=\mathcal{H}$ and
+    $\mathcal{L}^{*}\cup\mathcal{U}^{*}=\mathcal{H}$ and
     $\max_{h\in\mathcal{U}^{*}}\tfrac{M_{h}}{A_{h}}\leq\min_{h\in\mathcal{L}^{*}}\tfrac{m_{h}}{A_{h}}\mathrlap{\qquad\text{ if\, }\mathcal{U}^{*}\neq\emptyset\text{ and }\mathcal{L}^{*}\neq\emptyset,},$ (3.4)
@@
 In this section we introduce an algorithm solving Problem 1.1.
-In view of Theorem 3.1 its essential task is to split the set of all strata labels $\mathcal{H}$ into three subsets of take-min ($\mathcal{L}^{*}$), take-max ($\mathcal{U}^{*}$), and take-Neyman ((*$\mathcal{H}\setminus(\mathcal{L}^{*}\cup\mathcal{U}^{*})$).
+In view of Theorem 3.1 its essential task is to split the set of all strata labels $\mathcal{H}$ into three subsets of take-min ($\mathcal{L}^{*}$), take-max ($\mathcal{U}^{*}$), and take-Neyman ($\mathcal{H}\setminus(\mathcal{L}^{*}\cup\mathcal{U}^{*})$).
 We call this new algorithm RNABOX since it generalizes existing algorithm RNA in the sense that RNABOX solves optimum allocation problem with simultaneous lower and upper bounds, while the RNA is dedicated for the problem with upper bounds only, i.e.
@@
 Table 4.1: An example of RNABOX performance for a population with 10 strata and total sample size $n=5110$.
-Columns $\mathcal{L}_{r}/\mathcal{U}_{r}$, $r=1,\ldots,6$, represent the content of sets $\mathcal{L},\,\mathcal{U}$ respectively, in the $r$-th iteration of the RNABOX (between Step 4 and Step 5): symbols $\blacksquare$  or $\square$  indicate that a given stratum label $h$ is in $\mathcal{L}_{r}$ or $\mathcal{U}_{r}$, respectively.
-$h$
-$A_{h}$ $m_{h}$ $M_{h}$ $\mathcal{L}_{1}/\mathcal{U}_{1}$ $\mathcal{L}_{2}/\mathcal{U}_{2}$ $\mathcal{L}_{3}/\mathcal{U}_{3}$ $\mathcal{L}_{4}/\mathcal{U}_{4}$ $\mathcal{L}_{5}/\mathcal{U}_{5}$ $\mathcal{L}_{6}/\mathcal{U}_{6}$ $\mathbf{x}^{*}$
-1
-2700 750 900 $\square$ $\square$ $\square$ $\square$ $750$
-2
-2000 450 500 $\blacksquare$ $\square$ $\square$ $\square$ $\square$ $450$
-3
-4200 250 300 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $261.08$
-4
-4400 350 400 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $350$
-5
-3200 150 200 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $198.92$
-6
-6000 550 600 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $\square$ $550$
-7
-8400 650 700 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $650$
-8
-1900 50 100 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $100$
-9
-5400 850 900 $\blacksquare$ $\blacksquare$ $\square$ $\square$ $\square$ $850$
-10
-2000 950 1000 $\square$ $\square$ $\square$ $\square$ $\square$ $950$
+Columns $\mathcal{L}_{r}/\mathcal{U}_{r}$, $r=1,\ldots,6$, represent the content of sets $\mathcal{L},\,\mathcal{U}$ respectively, in the $r$-th iteration of the RNABOX (between Step 4 and Step 5): symbols $\square$  or $\blacksquare$  indicate that the stratum with label $h$ is in $\mathcal{L}_{r}$ or $\mathcal{U}_{r}$, respectively.
+$h$ $A_{h}$ $m_{h}$ $M_{h}$ $\mathcal{L}_{1}/\mathcal{U}_{1}$ $\mathcal{L}_{2}/\mathcal{U}_{2}$ $\mathcal{L}_{3}/\mathcal{U}_{3}$ $\mathcal{L}_{4}/\mathcal{U}_{4}$ $\mathcal{L}_{5}/\mathcal{U}_{5}$ $\mathcal{L}_{6}/\mathcal{U}_{6}$ $\mathbf{x}^{*}$
+1 2700 750 900 $\square$ $\square$ $\square$ $\square$ $750$
+2 2000 450 500 $\blacksquare$ $\square$ $\square$ $\square$ $\square$ $450$
+3 4200 250 300 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $261.08$
+4 4400 350 400 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $350$
+5 3200 150 200 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $198.92$
+6 6000 550 600 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $\square$ $550$
+7 8400 650 700 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\square$ $650$
+8 1900 50 100 $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $100$
+9 5400 850 900 $\blacksquare$ $\blacksquare$ $\square$ $\square$ $\square$ $850$
+10 2000 950 1000 $\square$ $\square$ $\square$ $\square$ $\square$ $950$
 SUM 5000 5600 0/8 1/7 3/6 4/3 5/3 7/1 $5110$
-For this example, RNABOX stops after $6$ iterations with take-min strata set *=$\mathcal{L}^{*}=\{1,2,4,6,7,9,10\}$, take-max strata set $\mathcal{U}^{*}=\{8\}$ and take-Neyman strata set (*$\mathcal{H}\setminus(\mathcal{L}^{*}\cup\mathcal{U}^{*})=\{3,5\}$ (see column $\mathcal{L}_{6}/\mathcal{U}_{6}$).
+For this example, RNABOX stops after $6$ iterations with take-min strata set =$\mathcal{L}^{*}=\{1,2,4,6,7,9,10\}$, take-max strata set $\mathcal{U}^{*}=\{8\}$ and take-Neyman strata set $\mathcal{H}\setminus(\mathcal{L}^{*}\cup\mathcal{U}^{*})=\{3,5\}$ (see column $\mathcal{L}_{6}/\mathcal{U}_{6}$).
 The optimum allocation is a regular one and it is given in column $\mathbf{x}^{*}$ of Table 4.1.
@@
 In view of Lemma C.2, using the notation introduced in Appendix C.1, in Step 3 of RNABOX, for $r^{*}\geq 2$ we have
-$*.\mathcal{U}_{r}=\{h\in\mathcal{H}\setminus\mathcal{L}_{r}:\,x^{**}_{h}=M_{h}\}\subseteq\mathcal{U}_{r-1},\mathrlap{\qquad r=2,\ldots,r^{*}.}$
+$\mathcal{U}_{r}=\{h\in\mathcal{H}\setminus\mathcal{L}_{r}:\,x^{**}_{h}=M_{h}\}\subseteq\mathcal{U}_{r-1},\mathrlap{\qquad r=2,\ldots,r^{*}.}$
 This suggests that the domain of discourse for $\mathcal{U}_{r}$ could be shrunk from $\mathcal{H}\setminus\mathcal{L}_{r}$ to $\mathcal{U}_{r-1}\subseteq\mathcal{H}\setminus\mathcal{L}_{r}$, i.e.
-$*.\mathcal{U}_{r}=\{h\in\mathcal{U}_{r-1}:\,x^{**}_{h}=M_{h}\},\mathrlap{\qquad r=2,\ldots,r^{*}.}$ (4.1)
-Given above observation and the fact that from the implementation point of view set $\mathcal{U}_{r}$ is determined internally by RNA, it is tempting to consider modification of RNA such that it makes use of the domain of discourse $\mathcal{U}_{r-1}$ for set $\mathcal{U}_{r}$.
+$\mathcal{U}_{r}=\{h\in\mathcal{U}_{r-1}:\,x^{**}_{h}=M_{h}\},\mathrlap{\qquad r=2,\ldots,r^{*}.}$ (4.1)
+Given the above observation and the fact that from the implementation point of view set $\mathcal{U}_{r}$ is determined internally by RNA, it is tempting to consider modification of RNA such that it makes use of the domain of discourse $\mathcal{U}_{r-1}$ for set $\mathcal{U}_{r}$.
 This domain could be specified as an additional input parameter, say $\mathcal{J}\subseteq\mathcal{H}$, and then Step 3 of RNA algorithm would read:
@@
 […]
-$h$
 […]
-1
-420 24 420 $30$ $54.44$
-2
-352 15 88 $88$blacksquare $45.63$
-3
-2689 1344 2689 $1344$square $1344$square
-4
-308 8 308 $22$ $39.93$
-5
-130 3 5 $5$blacksquare $5$blacksquare
-SUM
-1394 3510 $4891$/2/2 $4891$/1/3
-For this population, a naive version of the recursive Neyman stops in the second iteration and it gives allocation $\mathbf{x}$ with =$\mathcal{L}=\{3\}$ and $\mathcal{U}=\{2,5\}$, while the optimum choice is $\mathbf{x}^{*}$ with *=$\mathcal{L}^{*}=\{3\}$ and $\mathcal{U}^{*}=\{5\}$.
+1 420 24 420 $30$ $54.44$
+2 352 15 88 $88$blacksquare $45.63$
+3 2689 1344 2689 $1344$square $1344$square
+4 308 8 308 $22$ $39.93$
+5 130 3 5 $5$blacksquare $5$blacksquare
+SUM 1394 3510 $4891$/2/2 $4891$/1/3
+For this population, a naive version of the recursive Neyman stops in the second iteration and it gives allocation $\mathbf{x}$ with =$\mathcal{L}=\{3\}$ and $\mathcal{U}=\{2,5\}$, while the optimum choice is $\mathbf{x}^{*}$ with =$\mathcal{L}^{*}=\{3\}$ and $\mathcal{U}^{*}=\{5\}$.
 For non-optimum allocation $\mathbf{x}$, we have $s(\{3\},\,\{2,5\})=0.0714\ngeq 0.25=\tfrac{A_{2}}{M_{2}}$, i.e.
@@
 Numbers in brackets are the numbers of iterations of a given algorithm.
-In the case of RNABOX, it is a vector with number of iterations of the RNA (see Step 3 of RNABOX) for each outer iteration of RNABOX.
-Thus, the length of this vector is equal to the number of outer iterations of RNABOX.
+In the case of RNABOX, it is a vector with number of iterations of the RNA (see Step 3 of RNABOX) for each iteration of RNABOX.
+Thus, the length of this vector is equal to the number of iterations of RNABOX.
 Counts of take-min, take-Neyman, and take-max strata are shown on bottom graphs.
@@
 The gain in the execution time of the FPIA results from the fact that it typically runs through a smaller number of sets $\mathcal{L},\,\mathcal{U}\subset\mathcal{H}$ (according to (A.6)), than RNABOX in order to find the optimal $\mathcal{L}^{*}$ and $\mathcal{U}^{*}$.
-Although this approach usually gives correct results (as in the simulations reported in this section), it may happen that the FPIA misses the optimal sets *$\mathcal{L}^{*},\,\mathcal{U}^{*}\subsetneq\mathcal{H}$.
+Although this approach usually gives correct results (as in the simulations reported in this section), it may happen that the FPIA misses the optimal sets $\mathcal{L}^{*},\,\mathcal{U}^{*}\subsetneq\mathcal{H}$.
 In such a case, FPIA does not converge, as explained in Appendix A.2.
@@
 Consider the allocation for an example population as given in Table A.1.
-Table A.1: Two allocations for an example population with two strata: non-optimum $\mathbf{x}^{noptcond}$ with =$\mathcal{L}=\{1\},\,\mathcal{U}=\emptyset$, and optimum $\mathbf{x}^{*}$ with *$\mathcal{L}^{*}=\emptyset,\,\mathcal{U}^{*}=\{1\}$.
+Table A.1: Two allocations for an example population with two strata: non-optimum $\mathbf{x}^{noptcond}$ with =$\mathcal{L}=\{1\},\,\mathcal{U}=\emptyset$, and optimum $\mathbf{x}^{*}$ with $\mathcal{L}^{*}=\emptyset,\,\mathcal{U}^{*}=\{1\}$.
 Set function $s$ is defined in (3.1).
@@
 More specifically, Problem 1.1 is a convex optimization problem of a particular type in which inequality constraint functions (B.2) - (B.3) are affine.
-See the Appendix E for the definition of the convex optimization problem.
+See Appendix E for the definition of the convex optimization problem.
 Proof of Theorem 3.1.
 […]
+As explained in Remark B.1, Problem 1.1 is a convex optimization problem in which the inequality constraint functions $g_{h}^{m},\,g_{h}^{M},\,h\in\mathcal{H}$ are affine.
+The optimal solution for such a problem can be identified through the Karush-Kuhn-Tucker (KKT) conditions, in which case they are not only necessary but also sufficient; for further references, see Appendix E.
+The gradients of the objective function (B.1) and constraint functions (B.2) - (B.3) are as follows:
 $\nabla f(\mathbf{x})=(-\tfrac{A_{h}^{2}}{x_{h}^{2}},\,h\in\mathcal{H}),\quad\nabla w(\mathbf{x})=\underline{1},\quad\nabla g_{h}^{m}(\mathbf{x})=-\nabla g_{h}^{M}(\mathbf{x})=-\underline{1}_{h},\quad h\in\mathcal{H},$
@@
 Hence, the KKT conditions (E.2) for Problem 1.1 assume the form
-$-\tfrac{A_{h}^{2}}{x^{*2}_{h}}+\lambda-\mu_{h}^{m}+\mu_{h}^{M}=0,\mathrlap{\qquad h\in\mathcal{H},}$ (B.4)
+$-\tfrac{A_{h}^{2}}{{(x^{*}_{h})}^{2}}+\lambda-\mu_{h}^{m}+\mu_{h}^{M}=0,\mathrlap{\qquad h\in\mathcal{H},}$ (B.4)
 $\sum_{h\in\mathcal{H}}x_{h}^{*}-n=0,$ (B.5)
@@
 $\mu_{h}^{M}(x^{*}_{h}-M_{h})=0,\mathrlap{\qquad h\in\mathcal{H}.}$ (B.8)
-To prove Theorem 3.1, it suffices to show that for $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ with *$\mathcal{L}^{*},\,\mathcal{U}^{*}$ satisfying conditions of CASE I or CASE II, there exist $\lambda\in{\mathbb{R}}$ and $\mu^{m}_{h},\,\mu^{M}_{h}\geq 0,\,h\in\mathcal{H}$, such that (B.4) - (B.8) hold.
+To prove Theorem 3.1, it suffices to show that for $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ with $\mathcal{L}^{*},\,\mathcal{U}^{*}$ satisfying conditions of CASE I or CASE II, there exist $\lambda\in{\mathbb{R}}$ and $\mu^{m}_{h},\,\mu^{M}_{h}\geq 0,\,h\in\mathcal{H}$, such that (B.4) - (B.8) hold.
 It should also be noted that the requirement $m_{h}<M_{h},\,h\in\mathcal{H}$, guarantees that $\mathcal{L}^{*}$ and $\mathcal{U}^{*}$ defined in (3.3) and (3.4) are disjoint.
@@
   2. CASE II:
-    Take $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ with *$\mathcal{L}^{*},\,\mathcal{U}^{*}$ satisfying (3.4) and (3.5).
+    Take $\mathbf{x}^{*}=\mathbf{x}^{(\mathcal{L}^{*},\,\mathcal{U}^{*})}$ with $\mathcal{L}^{*},\,\mathcal{U}^{*}$ satisfying (3.4) and (3.5).
 Then, condition (B.5) becomes (3.5), while (B.6) is trivially met due to (3.2).
@@
 if $r^{*}\geq 2$), we get
-$*.\mathcal{L}_{r}=\mathcal{L}_{r-1}\cup\widetilde{\mathcal{L}}_{r-1}=\bigcup\limits_{i=1}^{r-1}\widetilde{\mathcal{L}}_{i},\mathrlap{\qquad r=2,\ldots,r^{*}.}$ (C.1)
+$\mathcal{L}_{r}=\mathcal{L}_{r-1}\cup\widetilde{\mathcal{L}}_{r-1}=\bigcup\limits_{i=1}^{r-1}\widetilde{\mathcal{L}}_{i},\mathrlap{\qquad r=2,\ldots,r^{*}.}$ (C.1)
 As RNABOX iterates, objects denoted by symbols $n$ and $\mathcal{H}$ are being modified.
@@
 In each iteration $r=1,\ldots,r^{*}\geq 1$, of RNABOX algorithm, a vector $(x^{**}_{h},\,h\in\mathcal{H}\setminus\mathcal{L}_{r})$ obtained in Step 3, has the elements of the form
-[…]
+$x^{**}_{h}=\begin{cases}M_{h},&\qquad{h\in\mathcal{U}_{r}\subseteq\mathcal{H}\setminus\mathcal{L}_{r}}\\ A_{h}\,s(\mathcal{L}_{r},\,\mathcal{U}_{r})<M_{h},&\mathrlap{\qquad h\in\mathcal{H}\setminus(\mathcal{L}_{r}\cup\mathcal{U}_{r}),}\end{cases}$ (C.12)
 where the set function $s$ is defined in (3.1). Equation (C.12) is a direct consequence of Theorem D.1.
@@
 $\mathcal{U}_{r^{*}}=\{h\in\mathcal{H}\setminus\mathcal{L}_{r^{*}}:\,A_{h}\,s(\mathcal{L}_{r^{*}},\,\mathcal{U}_{r^{*}})\geq M_{h}\},$ (C.14)
-if and only if $\mathbf{x}^{*}$ (computed at Step 6 of RNABOX algorithm) is: a regular allocation or a vertex allocation with $\mathcal{L}_{r^{*}}=\mathcal{H}$. Moreover, for $r^{*}\geq 1$,
-~$*.\widetilde{\mathcal{L}}_{r}=\{h\in\mathcal{H}\setminus(\mathcal{L}_{r}\cup\mathcal{U}_{r}):\,A_{h}\,s(\mathcal{L}_{r},\,\mathcal{U}_{r})\leq m_{h}\},\mathrlap{\qquad r=1,\ldots,r^{*}.}$ (C.15)
+if and only if $\mathbf{x}^{*}$ (computed at Step 6 of RNABOX algorithm) is: a regular allocation or a vertex allocation with $\mathcal{L}_{r^{*}}=\mathcal{H}$.
+Moreover, for $r^{*}\geq 1$,
+~$\widetilde{\mathcal{L}}_{r}=\{h\in\mathcal{H}\setminus(\mathcal{L}_{r}\cup\mathcal{U}_{r}):\,A_{h}\,s(\mathcal{L}_{r},\,\mathcal{U}_{r})\leq m_{h}\},\mathrlap{\qquad r=1,\ldots,r^{*}.}$ (C.15)
 Note that in Remark C.3, function $s$ is well-defined due to Remark C.1.
@@
 $s(\mathcal{L}_{r},\,\mathcal{U}_{r})\geq s(\mathcal{L}_{r+1},\,\mathcal{U}_{r+1}).$ (C.35)
-Hence (C.28) is proved. If $\mathbf{x}^{*}$ is a regular allocation, in view of Remark C.1, the same reasoning leading to inequality (C.35) clearly remains valid for $r=r^{*}-1,\,r^{*}\geq 2$. ∎
+Hence (C.28) is proved.
+If $\mathbf{x}^{*}$ is a regular allocation, in view of Remark C.1, the same reasoning leading to inequality (C.35) clearly remains valid for $r=r^{*}-1,\,r^{*}\geq 2$. ∎
 C.3 Proof of Theorem 4.1
@@
     $s(\mathcal{L}_{r^{*}-1},\,\mathcal{U}_{r^{*}-1})\leq\tfrac{m_{h}}{A_{h}},\mathrlap{\qquad h\in\mathcal{L}_{r^{*}-1}\cup\widetilde{\mathcal{L}}_{r^{*}-1}=\mathcal{L}_{r^{*}},}$ (C.46)
-    where the fact that above inequality is met for $h\in\mathcal{L}_{r^{*}-1}$ follows from (C.28).
+    where the fact that the above inequality is met for $h\in\mathcal{L}_{r^{*}-1}$ follows from (C.28).
 By comparing (C.45) and (C.46) we clearly see that (C.38) is satisfied.
@@
     $n-m_{\widetilde{\mathcal{L}}_{1}}-\ldots-m_{\widetilde{\mathcal{L}}_{r*-1}}=n-m_{\mathcal{L}_{r^{*}}}=M_{\mathcal{U}_{r^{*}}},$ (C.47)
-    where first equality follows from (C.1) while the second one follows from Step 3 of the RNABOX after referring to (D.3) and using the fact that $\mathcal{U}_{r^{*}}=\mathcal{H}\setminus\mathcal{L}_{r^{*}}$ for a vertex allocation.
+    where the first equality follows from (C.1) while the second one follows from Step 3 of the RNABOX after referring to (D.3) and using the fact that $\mathcal{U}_{r^{*}}=\mathcal{H}\setminus\mathcal{L}_{r^{*}}$ for a vertex allocation.
 Appendix D Appendix: Optimality conditions for Problem 2.1
@@
 Point $\mathbf{x}^{*}=(x_{h}^{*},\,h\in\mathcal{H})\in{\mathbb{R}}_{+}^{\lvert\mathcal{H}\rvert}$ is a solution to optimization Problem 2.1 if and only if $\mathbf{x}^{*}$ has entries of the form
-[…]
+$x_{h}^{*}=\begin{cases}M_{h},&\mathrlap{\qquad h\in\mathcal{U}^{*}}\\ A_{h}\,s(\emptyset,\,\mathcal{U}^{*}),&\mathrlap{\qquad h\in\mathcal{H}\setminus\mathcal{U}^{*},}\end{cases}$ (D.1)
 with $\mathcal{U}^{*}\subseteq\mathcal{H}$, such that one of the following two cases holds:
@@
 $\begin{split}\underset{\mathbf{x}\,\in\,\mathcal{D}}{\mathrm{minimize\leavevmode\nobreak\ \,}}&\quad f(\mathbf{x})\\ \mathrm{subject\leavevmode\nobreak\ to}&\quad w_{i}(\mathbf{x})=0,\quad{i=1,\ldots,k}\\ &\quad g_{j}(\mathbf{x})\leq 0,\quad{j=1,\ldots,\ell,}\end{split}$ (E.1)
-where $\mathcal{D}\subseteq{\mathbb{R}}^{p},\,p\in{\mathbb{N}}_{+}$, is the optimization variable, the objective function $f:\mathcal{D}_{f}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}}$ and inequality constraint functions $g_{j}:\mathcal{D}_{g_{j}}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}},\,j=1,\ldots,\ell$, are convex, whilst equality constraint functions $w_{i}:\mathcal{D}_{w_{i}}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}},\,i=1,\ldots,k$, are affine.
+where $\mathbf{x}$ is the optimization variable, $\mathcal{D}\subseteq{\mathbb{R}}^{p},\,p\in{\mathbb{N}}_{+}$, the objective function $f:\mathcal{D}_{f}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}}$ and inequality constraint functions $g_{j}:\mathcal{D}_{g_{j}}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}},\,j=1,\ldots,\ell$, are convex, whilst equality constraint functions $w_{i}:\mathcal{D}_{w_{i}}\subseteq{\mathbb{R}}^{p}\to{\mathbb{R}},\,i=1,\ldots,k$, are affine.
 Here, $\mathcal{D}=\mathcal{D}_{f}\cap\bigcap_{i=1}^{k}\mathcal{D}_{w_{i}}\cap\bigcap_{j=1}^{\ell}\mathcal{D}_{g_{j}}$ denotes a common domain of all the functions.
@@
 Theorem E.1 (KKT conditions for convex optimization problem with affine inequality constraints).
-A point $\mathbf{x}^{*}\in\mathcal{D}\subseteq{\mathbb{R}}^{p}$ is a solution to the convex optimization problem (E.1) in which functions $g_{j},\,j=1,\ldots,\ell$, are affine if and only if there exist numbers $\lambda_{i}\in{\mathbb{R}}$, $i=1,\ldots,k$, and $\mu_{j}\geq 0$, $j=1,\ldots,\ell$, called KKT multipliers, such that
-$\begin{gathered}\nabla f(\mathbf{x}^{*})+\sum_{i=1}^{k}\lambda_{i}\nabla w_{i}(\mathbf{x}^{*})+\sum_{j=1}^{\ell}\mu_{j}\nabla g_{j}(\mathbf{x}^{*})=\mathbf{0}\\ w_{i}(\mathbf{x}^{*})=0,\mathrlap{\qquad i=1,\ldots,k}\\ g_{j}(\mathbf{x}^{*})\leq 0\mathrlap{\qquad j=1,\ldots,\ell}\\ \mu_{j}g_{j}(\mathbf{x}^{*})=0\mathrlap{\qquad j=1,\ldots,\ell.}\end{gathered}$ (E.2)
+A point $\mathbf{x}^{*}\in\mathcal{D}\subseteq{\mathbb{R}}^{p},\,p\in{\mathbb{N}}_{+}$, is a solution to the convex optimization problem (E.1) in which functions $g_{j},\,j=1,\ldots,\ell$, are affine if and only if there exist numbers $\lambda_{i}\in{\mathbb{R}}$, $i=1,\ldots,k$, and $\mu_{j}\geq 0$, $j=1,\ldots,\ell$, called KKT multipliers, such that
+$\begin{gathered}\nabla f(\mathbf{x}^{*})+\sum_{i=1}^{k}\lambda_{i}\nabla w_{i}(\mathbf{x}^{*})+\sum_{j=1}^{\ell}\mu_{j}\nabla g_{j}(\mathbf{x}^{*})=\mathbf{0}\\ w_{i}(\mathbf{x}^{*})=0,\mathrlap{\qquad i=1,\ldots,k}\\ g_{j}(\mathbf{x}^{*})\leq 0,\mathrlap{\qquad j=1,\ldots,\ell}\\ \mu_{j}g_{j}(\mathbf{x}^{*})=0,\mathrlap{\qquad j=1,\ldots,\ell.}\end{gathered}$ (E.2)